### v1

# Safety-Driven Response Adaptive Randomisation: An Application in Non-inferiority Oncology Trials

Safety-Driven Response Adaptive Randomisation:
An Application in Non-inferiority Oncology Trials
[Uncaptioned image] Maria Vittoria  Chiaruttini
Unit of Biostatistics, Epidemiology and Public Health
University of Padova
Padova, Italy
mariavittoria.chiaruttini@studenti.unipd.it
&[Uncaptioned image] Lukas Pin
MRC Biostatistics Unit
University of Cambridge
Cambridge, UK
lukas.pin@mrc-bsu.cam.ac.uk
&[Uncaptioned image] Sofía S. Villar
MRC Biostatistics Unit
University of Cambridge
Cambridge, UK
sofia.villar@mrc-bsu.cam.ac.uk
(June 10, 2025)
The majority of response-adaptive randomisation (RAR) designs in the literature use efficacy data to dynamically allocate patients. Their applicability in settings where the efficacy measure is observable with a random delay, such as overall survival, remains challenging. This paper introduces a RAR design referred to as SAFER (Safety-Aware Flexible Elastic Randomisation) design , which uses early-emerging safety data to inform treatment allocation decisions in oncology trials. However, the design is applicable to a range of settings where it may be desirable to favour the arm demonstrating a superior safety profile. This is particularly relevant in non-inferiority trials, which aim to demonstrate an experimental treatment is not inferior to the standard of care, while offering advantages in terms of safety and tolerability. Consequently, an unavoidable and well-established trade-off arises for such designs: to balance the goals of preserving inferential efficiency for the primary non-inferiority outcome while incorporating safety considerations into the randomisation process through RAR. Our method, defines a randomisation procedure which prioritises the assignment of patients to better-tolerated arms and adjusts the allocation proportion according to the observed association between safety and efficacy endpoints. We illustrate our procedure through a comprehensive simulation study, inspired by the CAPP-IT Phase III oncology trial. Our results demonstrate that SAFER preserves statistical power even when efficacy and safety endpoints are weakly associated and offers power gains when a strong positive association is present. Moreover, the approach enables a faster/slower adaptation when efficacy and safety endpoints are temporally aligned/misaligned, respectively.
Keywords Project Optimus  $\cdot$ Patient-driven  $\cdot$ Quality Of Life  $\cdot$ Efficacy  $\cdot$ Adverse events  $\cdot$ Dose reduction  $\cdot$ Drug discontinuation  $\cdot$ Pharmacokinetic  $\cdot$ Pharmacodynamic  $\cdot$ Exposure–response
1 Introduction
In contemporary drug development, optimal dosing has become a central focus, aiming for a precise, patient-centred approach in oncology. Initiatives like the FDA’s Project Optimus [1] underscore this shift, emphasizing the best balance between drug’s efficacy and safety to maximize patient benefit and minimise toxicity. Traditionally, early-phase clinical trials primarily prioritise safety, often focusing on major adverse events, with efficacy being largely exploratory. As development progresses, later phases are driven by efficacy decisions, and while comprehensive safety data are collected, this information is not typically formally integrated into design decisions, for example through an adaptive design. This leads to a critical challenge: despite the goal of improving the overall benefit-risk profile and enabling patients to stay on treatment longer, late-phase designs often do not formally use accumulating safety data to adapt trial conduct to be better aligned with those aims.
The aforementioned issue becomes particularly relevant in the context of the modern oncology treatment landscape, which has been transformed by the advent of targeted therapies, monoclonal antibodies, and immune checkpoint inhibitors. These novel agents, designed to modulate specific molecular pathways, frequently demonstrate plateauing efficacy at doses well below the maximum tolerated dose (MTD) [2, 3]. Moreover, their administration typically spans prolonged treatment cycles and is often accompanied by chronic, low-grade adverse events (AEs), including fatigue, gastrointestinal disturbances, and dermatologic reactions. These AEs, although not classified as dose-limiting in the short windows of phase I trials, may significantly impair patient adherence, quality of life, and long-term tolerability in the later phases [4, 5]. A pertinent example to illustrate this is sotorasib (LUMAKRAS), a KRAS G12C inhibitor approved for non-small cell lung cancer [6]. For this treatment a 960 mg dose was selected for pivotal trials. Subsequent analyses showed that a 240 mg dose offered similar efficacy with a better safety profile, leading the FDA to mandate a post-marketing dose optimization study.
The above rationale motivates our interest in the following crucial question: how can a late phase adaptive design help identify alternative dosing strategies—such as lower or fractionated doses, different schedules or regiments—deliver, with comparable or superior efficacy with fewer side effects? This is particularly relevant to address the potential impact on phase II and III studies from dosing decisions from early trials often adopted on immediate and severe safety data. This is especially the case for agents with delayed efficacy signals and cumulative toxicity profiles that influence long-term adherence and outcomes. In light of the growing emphasis on toxicity management in oncology, we anticipate an increasing number of studies, both in early and late phase of development, will focus on evaluating alternative dosing regimens with improved safety profiles [7].
Our work aims to address the research question presented above by leveraging Response-Adaptive Randomisation (RAR). RAR refers to a class of statistical methodologies used in clinical trial design to dynamically adjust the probability of assigning patients to different treatment arms based on accumulating outcome data [8]. By increasingly favouring arms showing early superior performance, RAR can enhance both patient outcomes and the statistical efficiency of trials [9]. This adaptability is particularly advantageous when outcomes are available shortly after treatment initiation, allowing real-time modification of allocation proportions. However, in trials where primary efficacy outcomes are delayed – such as overall survival in oncology – RAR becomes more challenging due to the unavailability of timely, informative data during enrolment. Several strategies have been proposed to address this, including the use of early surrogate endpoints that correlate with long-term outcomes [10, 11, 12]. On the contrary, this paper introduces a RAR design that, alternatively, uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
However, a critical component of any safety-driven RAR is the inclusion on the possible interplay between safety and efficacy endpoints. While some drugs can maintain or improve efficacy after dose reductions, others could show reduced therapeutic benefit. From a statistical perspective, RAR designs may involve a trade-off with statistical power. In fact, while equal allocation does not necessarily maximise power in superiority trials [13], a balanced 1:1 allocation ratio is generally optimal in non-inferiority studies, where the alternative hypothesis assumes equal means (and variances) in the control and experimental arms. To this regards, Baldi Antognini et al. (2020) [14] introduced a compound optimal allocation strategy in survival endpoint trials to address the trade-off between "ethical" and inferential concerns in the choice of the randomisation proportion. Although their work is aligned with our rationale, it is limited to single-outcome efficacy trials since it does not capture the complexity of a joint safety-efficacy decisions.
Therefore, we propose an innovative Safety-Aware Flexible Elastic Randomisation (SAFER) design. SAFER dynamically adjusts patient allocation proportions based on accumulating safety endpoint data, aiming to strike a balance between patient safety, maximization of clinical benefit, and the exploitation of the association between efficacy and safety outcomes. This adaptive strategy is particularly suited for late-phase clinical trials, where it enhances decision-making while preserving statistical efficiency, especially with respect to study power.
The remainder of this paper is structured as follows. In Section 2, we define the key components of the SAFER design: (i) the efficacy endpoint, the associated power function, and sample size calculation; (ii) the definition of the safety endpoint; (iii) the target allocation proportion and the RAR procedure; and (iv) the integration of efficacy and safety data for the estimation of the allocation proportion. Section 3 introduces the CAPP-IT study as a motivating example and presents its re-design in a non-inferiority framework using the SAFER approach. Specifically, we outline the rationale and report the results from six distinct simulation scenarios, aimed at evaluating the performance of the proposed design under varying conditions. Section 4 discusses the simulation findings, explores broader implications of the SAFER design, and concludes with a summary of key contributions and directions for future research.
2 Components of the SAFER design
2.1 Efficacy Endpoint (Power and Sample Size calculation)
Our proposed design is specifically tailored to accommodate for the typical primary endpoint used in Phase II/III oncology trials. Specifically, we consider a time-to-event measure, such as Progression-Free Survival (PFS). Consequently, both the sample size and the primary analysis of our design will be expressed throughout this paper in terms of this designated endpoint. The sample size will be determined accordingly to ensure adequate statistical power for that main primary analysis. We will now formally outline this rationale, including the underlying assumptions and notation.
Let $j=E,C$ the experimental treatment and control arms, respectively. The patients’ primary efficacy endpoint (e) i.e., their survival time ($Y^{(e)}$ ) is assumed to be independent and identically distributed (i.i.d.) with probability density function (p.d.f) when treatment $j$ is assigned:
$f(y_{j}^{(e)};\theta_{j}^{(e)})=(\theta_{j}^{(e)-1})\exp\left(-\frac{y_{j}^{(e)}}{\theta_{j}^{(e)}}\right)$ (1)
where $\theta_{j}^{(e)-1}=\lambda_{j}^{(e)}\in\mathbb{R}^{+}$ denotes the event rate. Therefore, $\theta_{j}^{(e)}$ represents the mean of the treatment effect, and $\theta_{j}^{(e)2}$ represents the variance of the treatment effect. We will focus our design on the case in which the study will be testing for the non-inferiority of E with respect to C, the non-inferiority hypothesis for the mean difference becomes [15, 16]
$H_{0}:\theta_{C}^{(e)}-\theta_{E}^{(e)}\geq m\quad H_{1}:\theta_{C}^{(e)}-\theta_{E}^{(e)}<m,$ (2)
where $m$ in the non-inferiority margin. Based on this hypothesis, and under the assumptions that the means are strictly positive, and the standard regularity conditions hold ( to ensure the asymptotic normality of the Maximum Likelihood Estimations), we can construct a Wald statistic as follows [14]:
$W=\frac{\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)}-m}{\sqrt{Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})}}\sim\mathcal{N}(0,1),$ (3)
where the variance is estimated by
$Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})=\frac{1}{\sqrt{E}}\sqrt{\hat{\theta}_{C}^{(e)2}+\hat{\theta}_{E}^{(e)2}}$ (4)
We also incorporate an early stopping feature into our designs, recognizing that this feature, when combined with RAR, can offer advantages over RAR alone [17]. For simplicity, we have focused on only two analyses: an interim analysis (k=1) and the final analysis (k=2). Therefore, we will provide $W_{k=1}$ and $W_{k=2}$, estimated based on the first group available at the interim analysis and the complete sample at the end of follow-up, respectively.
We estimate the treatment effect through the Cox model [18] such that $W$ is approximated by $(\hat{\beta}_{E}-m)/SE(\hat{\beta}_{E})$ where $\hat{\beta}_{E}$ is the estimated difference between arms, i.e., $(\hat{\theta}_{C}^{(p)}-\hat{\theta}_{E}^{(p)})$ and $SE(\hat{\beta}_{E})$ is the square root of $Var(\hat{\beta}_{E})$. Although both the Cox model and exponential survival models are consistent estimators under exponentiality, in practice, the Cox model often converges more quickly and accurately in simulations. This is primarily due to its reliance on partial likelihood, fewer distributional assumptions, and better numerical conditioning in the presence of non-informative censoring and finite samples [19, 20].
Moreover, to estimate the minimum number of observed progression events $P$ needed for the study, we construct a $H_{0}$ rejection region as $R=\left\{-\infty,z_{1-\beta}\right\}$, with
$z_{1-\beta}=\sqrt{P\cdot(\log(HR_{H_{1}})-m)^{2}\pi_{E}\cdot(1-\pi_{E})}-z_{1-\alpha},$ (5)
where $\alpha$ is the significance level, $\beta$ is the Type II Error rate, $HR_{H_{1}}$ is the hazard ratio ($\lambda_{E}/\lambda_{C}$) under the alternative hypothesis, and $\pi_{E}$ is the realized allocation proportion to the experimental arm (E). We will let $\pi_{E}$ vary from 0.5 to 0.8. Therefore, for each $\pi_{E}$ and associated rejection region, we can calculate the minimum number of events $P$ needed to test the non-inferiority on the primary endpoint between treatments, based on the following power function definition [21].
[…]
After the calculation of the minimum number of events $P$ needed to achieve a study power of at least 0.8, we adjust the total number of events needed due to the application of the Lan-DeMets O’Brien-Fleming $\alpha$-spending function [22] to account for $k=2$ analyses. Finally, due to the survival primary endpoint, the overall sample size ($N$) is estimated by dividing the total number of events needed ($P$) by the probability to observe the event during the follow-up, given the event rate expected in the control and experimental arms, under the alternative hypothesis.
2.2 Safety Endpoint
The safety endpoint in our design is the time to dose-reduction/drug discontinuation due to low drug tolerability or AEs. The patient’s safety endpoint (s) ($Y^{(s)}$), is assumed to be i.i.d., with p.d.f. when treatment $j$ is assigned:
$f(y_{j}^{(s)};\theta_{j}^{(s)})=\theta_{j}^{(s)-1}\exp\left(-\frac{y_{j}^{(s)}}{\theta_{j}^{(s)}}\right)$ (7)
where $\theta_{j}^{(s)}$ is the average time to dose reduction/drug discontinuation in arm $j$. The safety model and its estimation will directly inform the RAR component of the design, as detailed in the next subsection.
2.3 Allocation Proportion Target
For our proposed design, the Neyman allocation [23, 24, 25] is a reasonable choice of allocation proportion to target. This is because Neyman allocation under these assumptions will consistently assign the majority of patients to the treatment that patients are observed to discontinue later [14] (or tolerate for the longest). The Neyman allocation proportion (or the proportion to the treatment arm $\pi_{E}$) can be estimated based on estimates of the safety average time to dose reduction as indicated below:
$\hat{\pi}_{E}=\frac{\hat{\theta}_{E}^{(s)}}{\hat{\theta}_{E}^{(s)}+\hat{\theta}_{C}^{(s)}}$ (8)
In our simulation design, we allow for $u=1,…,U$ update times for $\hat{\pi}_{E}$, which occur independently of the $k$-analyses defined by the group-sequential design (where interim efficacy analyses are assumed to occur only once, at $k=1$).
Specifically, the randomisation proportion $\hat{\pi}_{E}$ is updated sequentially:
  • For patients enrolled before the first update time ($u=1$), $\hat{\pi}_{E}$ is set to 0.5. This is the burn-in stage of the RAR feature of this design [26].
  • For patients enrolled between update time $u=1$ and $u=2$, $\hat{\pi}_{E}$ uses the estimate of $\hat{\theta}_{E}^{(s)}$ and $\hat{\theta}_{C}^{(s)}$ from the first update time.
  • This pattern continues, with patients enrolled between $u=2$ and $u=3$ using the estimate of $\hat{\theta}_{E}^{(s)}$ and $\hat{\theta}_{C}^{(s)}$ from the second update time, and so on, until the final update at $u=U$.
The number of update times ($U$) and the corresponding calendar occasions should be specified in the study protocol. The updates may be equally spaced across the enrolment period or not, depending on the design objectives. However, in the present design, we did not investigate how varying the number and timing of updates affects the design characteristics. Instead, we fixed equally spaced updates every three months, starting from month 3 through month 45, assuming a total enrolment period of 48 months. Crucially, each $\hat{\pi}_{E}$ is estimated using all available data up to a specific update time.
2.3.1 Elastic Adaptation of Allocation Proportion: Integration of Efficacy and Safety
A final element involves a function designed to modulate the RAR adaptation based on safety, utilizing the estimated association between safety and efficacy endpoints. The underlying idea is that if there’s sufficient evidence of non-inferiority, we would favour allocating more patients to the better-tolerated arms. Conversely, if non-inferiority is not supported by the current evidence in the study, the adaptation would shift away from those arms.
This is achieved by the inclusion in the design of an elastic function that dynamically adjusts the estimated allocation target proportion $\hat{\pi}_{E}$ based on the observed treatment effect on PFS, during the enrolment period. The SAFER function is defined in equation (9).
$SAFER(\hat{\pi}_{E})=\begin{cases}0.5&\text{if }\hat{\Phi}\leq 0.5\\ 0.5+(\hat{\pi}_{E}-0.5)\cdot\left(1-\left(1-\frac{\hat{\Phi}-0.5}{0.5}\right)^{\eta}\right)&\text{if }0.5<\hat{\Phi}<1\\ \hat{\pi}_{E}&\text{if }\hat{\Phi}=1\end{cases}$ (9)
Note that the maximum possible adjusted allocation target to be used in the design is equal to the standard adaptive allocation proportion $\hat{\pi}_{E}$ and the minimum adjusted allocation proportion is set to 0.5, representing equal allocation. The intermediate cases are introduced through a tuning parameter $\hat{\Phi}$ that quantifies the association between the safety and efficacy endpoints. Specifically, $\hat{\Phi}$ is estimated using the cumulative density function (CDF) of the absolute z-value, derived from a survival regression model, which corresponds to the estimated treatment effect on PFS. We only consider $\hat{\Phi}\geq 0.5$, effectively excluding the possibility of a negative association between treatment and PFS. Moreover, the parameter $\eta\geq 1$ is the elastic parameter which controls the curvature of the function: $\eta=1$ corresponds to a linear function, while $\eta>1$ results in a concave shape, as illustrated in Figure 1. Although convex shapes ($\eta<1$) are theoretically allowed, we did not explore the values of $\eta<1$, as the linear function is already considered a conservative choice to adjust the allocation proportion in cases of small $\hat{\Phi}$ values.
Figure 1: Target function of the SAFER design for two target allocations, $\hat{\pi}_{E}=0.7$ and $\hat{\pi}_{E}=0.9$, each evaluated under three shape parameters $\eta=1,2,5$. The SAFER function increases monotonically in $\Phi_{z}$ and reflects more aggressive allocation as both $\hat{\pi}_{E}$ and $\eta$ increase.
3 Performance Evaluation Through Simulation: redesigning a non-inferiority study
This section details the extensive simulation studies conducted to systematically assess the performance of the proposed SAFER design methodology. These simulations aim to evaluate key operating characteristics including statistical power and allocation patterns under various scenarios. The findings presented herein are based on a specific case study, which we redesign using SAFER and help us provide a comprehensive understanding of the SAFER’s design features and its potential advantages in different trial settings.
3.1 Motivating example: a follow-up to the CAPP-IT study
The motivating study is a multicentre, randomised, double-blind, placebo-controlled phase III trial included in the NIHR portfolio and conducted across 10 UK sites (CAPP-IT study; British Journal of Cancer, 2012, 107, 585–587 [27]). Patients with colorectal or breast cancer receiving palliative treatment with single-agent capecitabine were randomised to receive either concomitant pyridoxine or a matching placebo. The primary objective was to assess whether pyridoxine could reduce the need for capecitabine dose reduction and thereby improve patient outcomes. The safety endpoint was defined as the incidence of capecitabine dose reduction, while the efficacy endpoint was PFS. The original sample size calculation, based on the safety endpoint, required 270 patients to detect a reduction in dose reduction incidence from 30% to 15% with 80% power, accounting for potential dropouts. However, recruitment was slower than anticipated, and the study was closed prematurely after enrolling only 106 patients. Pyridoxine did not demonstrate a statistically significant effect on the objective response rate (odds ratio: 1.37; 95% CI: 0.475–3.96). Notably, there was a trend suggesting decreased PFS with pyridoxine: median PFS was 7.4 months in the pyridoxine group and 9.9 months in the placebo group (hazard ratio: 1.62; 95% CI: 0.91–2.88).
In the following subsections, we present a re-design of this study using our SAFER approach. Specifically, we formulate a non-inferiority hypothesis on the primary efficacy endpoint (PFS), while for the safety-driven RAR, we focus on the time to dose reduction or drug discontinuation due to capecitabine-induced side effects as the safety endpoint. Moreover, we assume a fixed total duration of the trial equal to 60 months, considering a uniformly distributed enrolment of 48 months and 12 months of follow-up. Assuming a monthly event rate ($\lambda_{C}$) of 0.069, which corresponds to a median PFS of approximately 10 months, and a $HR_{H0}$ of 1.25 ($\lambda_{E}/\lambda_{C}$), the median PFS in the experimental arm must exceed 8 months to demonstrate non-inferiority with respect to the control arm.
3.2 Modelling efficacy and safety endpoints association
To thoroughly assess SAFER’s performance, understanding the interplay between primary efficacy and safety/tolerability endpoints is essential. This section details how this relationship is incorporated into our framework, specifically addressing scenarios where these endpoints are not assumed to be independent.
[…]
Figure 2: Relationship between the ratio of average time-to-dose reduction/drug discontinuation and the ratio of average PFS in experimental (E) vs control (C) arms, at different level of endpoints association: independent, moderate, strong, very strong). Bootstrap (10,000 iterations) 95% confidence intervals.
3.3 Metrics to assess performance
Through simulations, we evaluated six distinct scenarios to quantify the trial’s performance under various conditions (described below), providing insights into its statistical properties and broader implications. For each scenario, we conducted 10,000 iterations, ensuring a Monte Carlo error of less than 0.5% for both power and type I error rate [28].
We report the overall study power, calculated as the proportion of simulated trials that rejected $H_{0}$ under the alternative hypothesis ($H_{1}$), aggregated across all simulation studies. Moreover, the estimated allocation proportion to the experimental arm is defined as the average proportion of patients assigned to treatment $E$ at the end of each simulated trial.
To assess the performance of the design and compare the SAFER approach with the fixed allocation proportion approach, we provide benchmark results—specifically, power and allocation proportions—obtained from a fixed randomisation design under various allocation proportions (from 0.5 to 0.8). These benchmarks were derived using a theoretical (non-simulation-based) approach, based on the normal approximation of the outcome as defined in equations (5) and (6), and are presented in Figure 3. The power loss is attributed to changes in the allocation proportion from 0.5 to 0.8. The overall power is maximized by $\pi_{E}=0.5$ since the $\lambda_{C}$ and $\lambda_{E}$ are assumed equal, and decreases to approximately 0.79 when $\pi_{E}=0.6$, 0.74 when $\pi_{E}=0.7$, and 0.63 when $\pi_{E}=0.8$.
Figure 3: Theoretical loss of power, using normal approximation, for the primary endpoint analysis corresponding to diverse Information Fractions of 50% and allocation proportions $\pi_{E}$, from 0.5 to 0.8.
Furthermore, to empirically verify the validity of the standard regularity conditions, we conducted a simulation using exponential distributions for survival outcomes, varying the information fraction (IF) at interim from 0.2 to 0.5; the results are reported in Table 1. The observed loss of power at increasing allocation proportions closely aligns with the theoretical results in Figure 3, suggesting that for further comparisons, the normal approximation can be reliably used in this setting. Interestingly, it seems that the IF does not modify the observed loss of power.
Table 1: Power using exponential distributed outcomes with fixed allocation proportion and varying information fraction at interim analysis.
{tblr}
cells = c, cell12 = c=4, vline2 = 1, vline2-5 = 2, hline1-3,10 = -, Allocation & Information Fraction
Fixed Proportion 0.2 0.3 0.4 0.5
0.5 0.80 0.80 0.80 0.80
0.55 0.79 0.79 0.79 0.79
0.6 0.79 0.79 0.78 0.78
0.65 0.76 0.76 0.76 0.75
0.7 0.74 0.74 0.74 0.73
0.75 0.69 0.69 0.69 0.70
0.8 0.64 0.63 0.63 0.63
3.4 Simulating scenarios
In this section, we will describe in detail the six distinct scenarios used for the simulation study.
Scenario 1: Safety-driven RAR design without a SAFER approach: impact on power under independence between the efficacy and safety endpoints.
This scenario assesses the impact of a safety-driven RAR design—without implementing the SAFER approach—on power, under the assumption of independence between the efficacy and safety endpoints. Theoretical results (in section 3.3) suggest that this impact is expected to be negative, implying a loss in power. Given that we set $U=15$, $\hat{\pi}_{E}$ is updated every three months until month 45, following an initial burn-in period during which a balanced randomisation is applied for the first three months. The median time to dose reduction or drug discontinuation is simulated at 1.5 months in the control group, while in the experimental group, it is varied across values of 1.5, 1.8, 2.25, 2.8, 3.5, 4.5, and 6 months to achieve target allocation proportions (based on the Neyman allocation) of 0.5, 0.55, 0.60, 0.65, 0.70, 0.75, and 0.80, respectively.
Scenario 2: Safety-driven RAR design with a SAFER approach: impact on power under diverse level of dependence between the efficacy and safety endpoints.
This scenario is designed as a counterpart to scenario 1 and aims to assess the impact of changing the allocation proportion on statistical power when the SAFER design is implemented. Specifically, we simulated increasing levels of association between the endpoints, with $E[\gamma_{1}]$ set to 0.001, 0.005, 0.01, 0.03, and 0.05, corresponding to very weak, weak, moderate, strong, and very strong associations, respectively. Under this scenario, we expect power to be preserved in the presence of weak dependence between endpoints, and to increase as dependence strengthens, potentially due to the enhanced effectiveness of the experimental treatment on the efficacy endpoint driven by prolonged well-tolerated drug exposure.
Scenario 3: Varying elasticity of the target function in the SAFER design.
This scenario is designed to evaluate the impact of the elastic parameter $\eta$ on the SAFER approch results. Specifically, scenarios 3a and 3b report the results of simulations conducted with $\eta=1$ and $\eta=5$, respectively. These values correspond to the absence of elasticity and a high degree of elasticity in the target SAFER function, as reported in Figure 1.
Scenario 4: Differential timing in the observation of endpoint data.
This scenario aims to assess the speed at which the allocation proportion reaches the target level, when the median times of the safety and efficacy endpoints are either close or far apart i.e., delayed efficacy outcome [11]. For instance, we expect a faster pattern when the median PFS is close to the median time to dose reduction or drug discontinuation because there will be enough PFS events for the $\Phi$ parameter estimation, early in the recruitment period. On the other hand, when evidence of the positive association between safety and efficacy emerges later in the recruitment, we expect the SAFER function to down-weight the allocation proportion for a longer period until sufficient evidence on PFS becomes available. Under this scenario, the median PFS ranges from 3 to 24 months.
Scenario 5: Impact of informative drop-out (related to the safety endpoint).
This scenario aims to assess the impact of informative drop-out on the operating characteristics of the proposed design. We define informative drop-out as a situation in which a patient is observed until experiencing a side effect and subsequently drops out of the study. In practice, this means that the patient is followed until dose reduction or drug discontinuation occurs, after which neither the time to PFS nor the occurrence of the progression event is observed. We simulated the average allocation proportion and power under varying informative drop-out rates, ranging from 5% (very low) to 25% (very high), with drop-outs uniformly distributed over the accrual period. To address the potential bias introduced by informative drop-out, we considered the strategies for handling intercurrent events as defined in the Estimands framework [29]. Specifically, we considered the composite strategy, where the time-to-PFS is defined as the last available follow-up time, with the occurrence of the progression = yes.
Scenario 6: Impact of under-reported safety events.
This scenario aims to assess the impact of under-reported safety events on the operating characteristics of the proposed design. Unlike the informative drop-out phenomenon—where patients are observed until an adverse event occurs and then drop out—in the case of under-reported safety events, it is not recorded at all and only the survival outcome (i.e., progression and the corresponding time-to-PFS) is available [30].
3.5 Simulation Results
The total sample size required to achieve at least a power value of 0.8 is 881 (about 18 patients enrolled per months). However, due to the interim analysis, the total number of events needed increases to 888 when IF = 0.5.
Table 2 presents the results of the simulations under scenario 1 (independence between endpoints). Both the estimated allocation proportion and the observed loss of power closely align with the target allocation and the theoretical power, respectively, with only a minor discrepancy with respect to Table 1. This slight difference can be attributed to the burn-in period, since the results in Table 1 are derived using the unequal allocation proportion without any burn-in period, while scenario 1 includes this initial adjustment phase.
Table 2: Estimated allocation proportion and power summary: Scenario 1
{tblr}
cells = c, cell11 = c=2, cell13 = c=4, vline2 = 1, vline2-6 = 2, vline3 = 3-9, hline1,10 = -0.08em, hline2-3 = -, Allocation & Information Fraction
Target Estimated 0.2 0.3 0.4 0.5
0.5 0.5 0.80 0.80 0.80 0.80
0.55 0.54 0.80 0.79 0.79 0.79
0.6 0.59 0.78 0.79 0.78 0.78
0.65 0.64 0.77 0.77 0.76 0.76
0.7 0.69 0.74 0.74 0.73 0.73
0.75 0.73 0.71 0.71 0.70 0.70
0.8 0.78 0.66 0.66 0.65 0.66
Table 3 presents the results of the simulations under scenario 2, 3a, and 3b. In scenario 2, the estimated allocation proportion closely follows the target allocation, reflecting the true difference between treatment arms in terms of time to dose reduction or drug discontinuation. When there is no difference between arms on the safety endpoint (i.e., target allocation = 0.5), the power remains stable at approximately 0.8. When $E[\gamma_{1}]$ ranges from 0.001 to 0.005 (very weak or weak association between endpoints), the increasing target allocation leads to a loss of power in the efficacy endpoint analysis. On the contrary, when $E[\gamma_{1}]$ increases, reaching 0.03 and 0.05 values (strong or very strong positive association between endpoints), we observe a power gain. This improvement is driven by a biologically meaningful positive effect on PFS due to a drug exposure free from AEs (expressed by the model in section 3.2).
Under scenario 3a, the SAFER approach is implemented with the elastic parameter $\eta$ fixed at 1. In the absence of endpoints association, the SAFER approach effectively shrinks the estimated allocation towards 0.5, thereby preserving statistical power. Unlike scenario 2, where power losses were observed, when the endpoints association is very weak or weak, in scenario 3a the power of the efficacy analysis consistently remains above 0.8, regardless of the increase in target allocations. Moreover, as expected, higher allocation proportions are observed for stronger association between endpoints and more extreme target allocations. The maximum allocation proportion achieved in scenario 3a is approximately 0.7. This reflects the effect of setting $\eta=1$, which enforces strong regularization and slows the adaptation of allocation proportions.
On the other hand, in scenario 3b, when $\eta=5$, the average allocation reaches 0.74 when the target is 0.8. These results empirically demonstrate the role of $\eta$: the higher its value, the closer the estimated average allocation approaches the target allocation. Furthermore, regardless of the elasticity of the target function, when the relationship with the safety endpoint is weak, the SAFER approach consistently shrinks $\hat{\pi}_{E}$ downward. This mechanism minimises the potential loss of power in the efficacy analysis.
Table 3: Estimated allocation proportion and power summary: Scenarios 2, 3a, and 3b (IF = 0.5)
Target allocation Endpoint association Scenario 2 Scenario 3a ($\eta=1$) Scenario 3b ($\eta=5$)
$\hat{\pi}_{E}$ (P) $SAFER(\hat{\pi}_{E})$ (P) $SAFER(\hat{\pi}_{E})$ (P)
0.5 Very Weak 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
Weak 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
Moderate 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
Strong 0.50 (0.79) 0.50 (0.79) 0.50 (0.79)
Very Strong 0.50 (0.79) 0.50 (0.79) 0.50 (0.79)
0.6 Very Weak 0.59 (0.78) 0.52 (0.80) 0.54 (0.80)
Weak 0.59 (0.80) 0.53 (0.82) 0.54 (0.82)
Moderate 0.59 (0.81) 0.53 (0.83) 0.54 (0.83)
Strong 0.60 (0.86) 0.53 (0.87) 0.55 (0.87)
Very Strong 0.60 (0.89) 0.53 (0.91) 0.55 (0.91)
0.7 Very Weak 0.69 (0.74) 0.55 (0.81) 0.58 (0.81)
Weak 0.69 (0.78) 0.55 (0.84) 0.59 (0.84)
Moderate 0.69 (0.81) 0.56 (0.87) 0.59 (0.87)
Strong 0.69 (0.92) 0.58 (0.95) 0.62 (0.95)
Very Strong 0.69 (0.96) 0.60 (0.98) 0.64 (0.99)
0.8 Very Weak 0.78 (0.67) 0.57 (0.81) 0.62 (0.81)
Weak 0.78 (0.73) 0.59 (0.87) 0.64 (0.86)
Moderate 0.78 (0.79) 0.60 (0.92) 0.66 (0.92)
Strong 0.79 (0.94) 0.66 (0.99) 0.71 (0.99)
Very Strong 0.79 (0.99) 0.70 (1.00) 0.74 (1.00)
P = Power; $\hat{\pi}_{E}$ = estimated allocation using Neyman approach; $SAFER(\hat{\pi}_{E})$ = estimated allocation using SAFER approach.
Figure 4 and Figure 5 illustrate the median allocation proportion, estimated using the SAFER design, over the accrual period of 48 months. Scenario 4 is specifically designed to evaluate the differences in the speed at which the target allocation proportion is reached. This is assessed under conditions where the median times to the safety and efficacy endpoints are either close together or more distant. As expected, a faster pattern of reaching the target is observed when the median PFS is close (3 months) to the median time to dose reduction or drug discontinuation. Notably, when the median PFS is 3 months, a significant reduction in variability of allocation proportion estimation is observed already from the 6th month (Figure 4). Moreover, when the difference in safety outcome between arm decreases i.e., the target allocation is 0.7 (Figure 5), the reduction in variability of allocation proportion estimation, in the panel A, is observed from the 12th month. Moreover, as the median PFS increases (panel B-D), the evidence of the positive association between the endpoints becomes more apparent later in the recruitment period. As a result, the SAFER elastic function consistently slows down the approach to the target allocation when efficacy and safety endpoints are temporally misaligned.
Figure 4: Median and IQR of allocation proportions under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;). Target allocation=0.8. IF=0.5. $\eta$=5. $E[\gamma_{1}]$=0.05
Figure 5: Median and IQR of allocation proportions under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;). Target allocation=0.7. IF=0.5. $\eta$=5. $E[\gamma_{1}]$=0.05
Table 4 reports the average allocation proportion and the estimated power at increasing percentages of drop-out and under-reporting rates. A direct comparison with the values reported in Table 2 corresponding to scenario 3b and association "Very Strong" is possible since both $E[\gamma_{1}]$ and $\eta$ are fixed at the same levels.
In scenario 5, as the drop-out rate increases, the allocation proportion also increases. For instance, when the target allocation is 0.8, it rises from 74% (no drop-out in Table 2) to 77% with a drop-out rate of 25%. Interestingly, the estimated power also increases with higher drop-out rates. Although this result might seem counter-intuitive, a detailed explanation is provided in the discussion section.
On the other hand, in scenario 6, as the under-report rate increases, the allocation proportion decreases. This result is expected, as the core of the proposed RAR method relies on detecting differences in the safety endpoint. The greater the under-reporting of this endpoint, the less evidence the algorithm has to adjust the allocation proportion in favour of the better-tolerated arm.
Table 4: Estimated allocation proportion and power summary: Scenarios 5 and 6 (IF=0.5, $E[\gamma]=0.05$, $\eta=5$)
Target allocation Rate level Scenario 5 (Drop-out) Scenario 6 (Under-reporting)
$SAFER(\hat{\pi}_{E})$ (P) $SAFER(\hat{\pi}_{E})$ (P)
0.5 Very Low 0.50 (0.81) 0.50 (0.79)
Low 0.50 (0.81) 0.50 (0.79)
Moderate 0.50 (0.83) 0.50 (0.79)
High 0.50 (0.85) 0.50 (0.79)
Very High 0.50 (0.87) 0.50 (0.79)
0.6 Very Low 0.55 (0.92) 0.55 (0.91)
Low 0.56 (0.94) 0.54 (0.91)
Moderate 0.56 (0.95) 0.54 (0.91)
High 0.56 (0.96) 0.54 (0.90)
Very High 0.56 (0.97) 0.53 (0.91)
0.7 Very Low 0.64 (0.99) 0.63 (0.98)
Low 0.65 (0.99) 0.62 (0.98)
Moderate 0.65 (1.00) 0.61 (0.99)
High 0.66 (1.00) 0.60 (0.99)
Very High 0.67 (1.00) 0.60 (0.98)
0.8 Very Low 0.75 (1.00) 0.73 (1.00)
Low 0.76 (1.00) 0.72 (1.00)
Moderate 0.76 (1.00) 0.71 (1.00)
High 0.77 (1.00) 0.70 (1.00)
Very High 0.77 (1.00) 0.69 (1.00)
P = Power; $SAFER(\hat{\pi}_{E})$ = Estimated allocation using SAFER approach.
4 Discussion
This paper introduces a safety-driven RAR– namely SAFER design – tailored for non-inferiority trials which aims to identify alternative curative strategies—such as lower or fractionated doses, different schedules, regiments—deliver, or combo therapies with comparable efficacy to the standard of care but higher tolerability and fewer side effects. By incorporating safety data into the randomisation process, our method enhances patient-oriented research by preferentially allocating participants to treatment arms with a more favourable adverse event or side effect profile, while preserving statistical power. This approach is particularly relevant in oncology, where novel agents—often targeting specific molecular pathways—are typically administered over prolonged treatment cycles and are frequently associated with chronic toxicities, which constitute a major cause of treatment discontinuation.
Main Results
Our results confirm the theoretical expectation that unbalanced allocation in two-arm non-inferiority trials leads to a power loss in the efficacy endpoint analysis. The use of response-adaptive designs in the context of non-inferiority has been previously explored in the methodological works of [31, 32], and applied in clinical trials by [33, 34]. These studies, however, define the study power in Bayesian perspective for multi-arm clinical trials, where the goal is to correctly identify as many non-inferior arms as possible among multiple experimental groups, under the assumption that both inferior and non-inferior treatments exist relative to a common control. Such approaches are not directly transferable to the more common two-arm non-inferiority trials, particularly prevalent in phase III oncology studies. In these settings, the alternative hypothesis explicitly posits that the experimental treatment is non-inferior to the control, and the study power is defined as the probability of correctly rejecting the null hypothesis of inferiority—i.e., demonstrating that the estimated treatment effect lies within the pre-specified non-inferiority margin.
The trade-off between safety-driven allocation proportion and statistical power is particularly relevant in the context of oncology, where assigning more patients to better-tolerated arms mat improve patient experience and treatment effectiveness. The proposed SAFER approach offers a practical solution by dynamically adjusting the allocation proportion based on the observed association between safety and efficacy. When no association is observed, the SAFER approach appropriately shrinks the allocation back toward complete randomisation (1:1), preserving statistical power. When a strong (weak) positive association emerges, the algorithm allows for more (less) aggressive adaptation without undermining statistical validity. Moreover, the elasticity parameter ($\eta$) can regulate the speed and extent of adaptation. Smaller values of $\eta$ impose greater regularization, making the allocation proportion more conservative (close to 0.5), whereas higher values enable more aggressive deviation from balanced allocation.
Moreover, we underscore the critical role of endpoint timing in the proposed design (scenario 4). When the safety and efficacy endpoints are temporally aligned, the design detects associations more rapidly, enabling faster and more confident adjustments in allocation proportions. In contrast, greater temporal separation between endpoints delays the detection of such associations and slows the adaptation process. This finding highlights a key advantage of the SAFER framework: the flexibility provided by the proposed function allows for better control over the degree of adaptation, particularly in settings with delayed efficacy outcomes—a common feature in survival studies. Unlike Sequential Maximum Likelihood Estimation RAR approach, which may adjust allocations based solely on safety endpoint data, the SAFER design incorporates regularization that prevents premature and extreme imbalances in allocation. This is especially relevant in non-inferiority settings, where the experimental arm could ultimately prove inferior. By moderating allocation shifts until sufficient evidence accumulates on the efficacy endpoint, the design protects patients from being preferentially assigned to an arm that performs better on safety outcomes but may fail on efficacy assessment. This guarantees that, under the null hypothesis, the allocation mechanism does not favour an inferior treatment, even if it appears more tolerable.
Further, the study assessed the impact of two common issues in oncology trials: informative drop-out and under-reported adverse events. The adoption of a composite strategy to address informative drop-out results in an increase in the estimated allocation proportion as the drop-out rate rises (Table 4). Specifically, under this strategy, the time-to PFS is imputed as equal to the time-to-drug reduction for patients who discontinue prematurely. Consequently, the allocation proportion gradually converges toward $\hat{\pi}_{E}$, regardless the adaptation of the SAFER approach. This convergence occurs because the observed differences in PFS between treatment arms become increasingly similar to the differences in time-to-drug reduction, thereby diminishing the influence of the down-weighting effect introduced by the SAFER approach. In parallel, statistical power also increases. This is attributable to the growing number of events contributing to the composite endpoint (safety plus efficacy) as the drop-out rate rises. Additionally, by design, the treatment arms show a substantial difference on the safety endpoint, which enlarge the overall effect size captured by the composite measure. As a result, the likelihood of rejecting the null hypothesis increases.
While a higher drop-out rate introduces bias into the estimation process, under the alternative hypothesis, this bias may not necessarily be detrimental to patient outcomes. However, a problematic scenario may arise if the experimental arm demonstrates better tolerability but worse efficacy. In such cases, the composite strategy may fail to protect the efficacy endpoint analysis. Specifically, imputing the time-to-PFS as equal to the time of dose reduction or discontinuation may erroneously lead to the rejection of the null hypothesis, thereby substantially increasing the risk of Type I error rate inflation. Importantly, this bias is not introduced by the RAR design itself, but rather by the nature of the composite strategy—particularly when drop-out is more likely in the presence of poor tolerability, and tolerability differs significantly between arms. Therefore, when adopting the composite strategy, the treatment effect on the efficacy endpoint should be interpreted as a combined measure of both efficacy and safety since it reflects the integrated impact of tolerability on clinical outcomes. Under this interpretation only, the use of a composite strategy is reasonable and justified.
As an alternative, when a biologically plausible and well-specified association exists between the efficacy and safety endpoints, a hypothetical strategy may be used. Within this framework, the occurrence of the efficacy endpoint can be modeled conditionally on the observed safety outcome, thereby estimating what would have occurred in the absence of the intercurrent event [35]. This approach can preserve estimand consistency and potentially improve inference efficiency, provided the assumptions underlying the association model are adequately met. On the other hand, if the efficacy data are available even for patients who drop out, a treatment policy strategy could be adopted. This would mitigate the risk of biased inference due to informative censoring [29, 35]. In general, further analyses and comprehensive simulation studies are necessary to thoroughly assess the effects of informative drop-out under varying scenarios and parameter settings and the best way to handle the bias in this setting.
Considerations on design flexibility and generalizability
The current design adopts a group-sequential framework with a single interim analysis and a final analysis. However, it is fully customizable, allowing for the specification of multiple interim analyses and the selection of various $\alpha$-spending functions. Additionally, a $\beta$-spending function can be incorporated to formally account for futility stopping, thereby providing greater flexibility in controlling both Type I and Type II error rates [36].
Although the proposed design is developed within a frequentist framework, a Bayesian counterpart can be readily implemented. Specifically, the allocation proportion parameter $\hat{\pi}_{E}$ could be estimated using Thompson Sampling or derived from posterior probability-based decision rules [37]. Similarly, the parameter $\Phi$ can be interpreted as the posterior probability that the treatment effect exceeds zero. This Bayesian extension preserves the adaptive nature of the design while offering enhanced flexibility to incorporate prior information and dynamically update beliefs as data accumulate.
An important methodological consideration pertains to the generalizability of the SAFER design. While the motivating example is situated within a non-inferiority framework, the design is equally applicable to superiority trials. A central message of this work is to highlight the role of safety data as a guiding component in RAR designs, owing to their early availability and their potential to enhance patient-centred research. Furthermore, both the $\arg\min(\hat{\pi}_{E}^{\text{SAFER}})$ and $\arg\max(\hat{\pi}_{E}^{\text{SAFER}})$, as well as the function boundaries (which, in our example, are defined as $\text{lower}=0.5$ and $\text{upper}=1$), can be tailored to the specific requirements of the trial setting.
Furthermore, a transition from non-inferiority to superiority testing could be even considered within the same trial framework, especially in scenarios where safety and efficacy endpoints are positively associated. Our simulation results demonstrate that the design, simulated under positive association between safety and efficacy, achieves power exceeding the nominal 80% planned during sample size calculation. This surplus is attributable to the conservative nature of non-inferiority testing, wherein the pre-specified margin inflates the rejection region and increases the probability of rejecting the null hypothesis when the experimental treatment outperforms expectations. Formally, superiority can be concluded if the 95% confidence interval for the treatment effect not only excludes the non-inferiority margin but also lies entirely above zero, thereby rejecting the null hypothesis of no difference at the conventional $\alpha$ level [38]. By reformulating the hypothesis test to directly compare mean treatment effects—i.e., $H_{0}:\theta_{E}\leq\theta_{C}$ versus $H_{1}:\theta_{E}>\theta_{C}$—and removing the non-inferiority margin from the test statistic, the same dataset can be reanalyzed for superiority. Notably, this procedure leverages existing data without requiring an increase in sample size. In favourable scenarios of positive association between endpoints, the observed power surplus could thus be reallocated to formally test for superiority, enhancing the trial’s inferential strength and its potential clinical relevance.
Considerations on association between endpoints
The proposed design, for simulation purposes, assumes a functional relationship between the cumulative number of completed drug cycles and the PFS rate, under the hypothesis that dose reductions/drug discontinuation adversely impact treatment efficacy. Specifically, the design models this association to reflect the plausible clinical scenario where reduced exposure to full-dose treatment diminishes therapeutic benefit. Although the motivating case study—the CAPP-IT trial—compares standard treatment versus standard treatment plus pyridoxine (experimental arm), the proposed methodological framework is generalizable to alternative settings. For instance, it can be applied to trials comparing two distinct treatment regimens [39], schedules [40], or dosing strategies [41], whether in monotherapy or combination settings. Adaptation to such contexts would simply require modification of the underlying association model linking the safety and efficacy endpoints.
Importantly, it should be noted that the specification of the association model between safety and efficacy endpoints is required exclusively for simulation purposes to assess the operating characteristics of the SAFER design. In the actual trial implementation, no explicit modeling of this relationship is necessary to drive allocation decisions. The estimation of the $\Phi$ parameter—quantifying the empirical strength of the association—is derived directly from the same generalized linear model specified for the efficacy endpoint analysis, formally defined as $Y^{(p)}\sim j\in{E,C}$, where $E$ and $C$ denote the experimental and control arms, respectively. This model serves a dual purpose within the proposed adaptive framework. First, it is applied periodically—for instance every three months—to update the allocation proportions, where it functions solely to compute the test statistic ($z$-value) and its corresponding C.D.F. This C.D.F. is then used to inform the SAFER allocation mechanism without performing any interim hypothesis testing or inferential decision-making, avoiding uncontrolled multiple testing issues [42]. Second, the same model underpins the formal inferential framework at both interim and final analyses, where the computed $z$-value is evaluated against pre-specified decision boundaries to support statistical testing of the primary hypothesis.
Furthermore, it is worth highlighting that the clinical rationale for favouring allocation towards the arm demonstrating superior performance on the safety endpoint remains justifiable, even under the assumption of independence between endpoints and the associated loss of statistical power. Since the loss of power under varying allocation proportions is quantifiable, trial designs could, resources permitting, incorporate a pre-specified sample size adaptation rule at interim analyses. This would allow for the recovery of power losses while preserving the responsiveness of the randomisation mechanism [43].
Considerations on safety endpoint analysis
Although the present work is centred on hypothesis testing for the efficacy endpoint, the proposed design framework is sufficiently flexible to incorporate safety endpoints without loss of generality. These endpoints can not only be leveraged to inform the randomisation algorithm, but may also be elevated to co-primary status. In such cases, appropriate control of the family-wise Type I error rate can be achieved through standard procedures (e.g., Bonferroni correction, Holm’s method) [44] or more advanced alpha-spending strategies and multiplicity adjustment techniques tailored for hierarchical endpoints [45].
Limitations of the parametric assumptions
Although the exponential distribution is commonly used for simulating survival trial designs, it introduces limitations due to the parametric assumptions it imposes. However, the SAFER design is highly customizable: both the distributional assumptions and the type of endpoint can be relaxed. In the current work, we adopted time-to-event endpoints, but without loss of generality, the proposed design can also accommodate other commonly used endpoints, such as binary or count outcomes, under both parametric and non-parametric frameworks.
Conclusions
Ultimately, our proposed safety-driven RAR design offers a significant contribution in terms of practical tools to deliver patient-centred trials by adaptive learning from quickly observable safety data to proactively mitigate the assignment to poorly tolerated treatments without negatively affecting the power of the primary trial’s objective. While highly relevant to Phase III oncology, its principles are broadly applicable. Future efforts will focus on extending this framework to multi-arm trials and incorporating richer safety outcome data, including digital endpoints for more frequent adaptation. Methodological research will explore alternative distributions and non-parametric methods for small samples, with empirical validation in oncology settings essential for practical refinement.
Acknowledgments
The authors acknowledge the use of large language models (LLMs) to assist with generating figures and tables and refining grammar and wording in this paper. The LLMs were not used for data analysis, interpretation, or original scientific writing. All content has been carefully reviewed and verified by the authors, who take full responsibility for the integrity and accuracy of the work. Moreover, the authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP), the Cusanuswerk e.V. (LP), and the PhD scholarship from the University of Padua (MVC). SSV is part of PhaseV’s advisory board.

### v2

@@
 sofia.villar@mrc-bsu.cam.ac.uk
-(June 10, 2025)
-The majority of response-adaptive randomisation (RAR) designs in the literature use efficacy data to dynamically allocate patients.
-Their applicability in settings where the efficacy measure is observable with a random delay, such as overall survival, remains challenging.
-This paper introduces a RAR design referred to as SAFER (Safety-Aware Flexible Elastic Randomisation) design , which uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
-However, the design is applicable to a range of settings where it may be desirable to favour the arm demonstrating a superior safety profile.
-This is particularly relevant in non-inferiority trials, which aim to demonstrate an experimental treatment is not inferior to the standard of care, while offering advantages in terms of safety and tolerability.
-Consequently, an unavoidable and well-established trade-off arises for such designs: to balance the goals of preserving inferential efficiency for the primary non-inferiority outcome while incorporating safety considerations into the randomisation process through RAR.
-Our method, defines a randomisation procedure which prioritises the assignment of patients to better-tolerated arms and adjusts the allocation proportion according to the observed association between safety and efficacy endpoints.
-We illustrate our procedure through a comprehensive simulation study, inspired by the CAPP-IT Phase III oncology trial.
-Our results demonstrate that SAFER preserves statistical power even when efficacy and safety endpoints are weakly associated and offers power gains when a strong positive association is present.
-Moreover, the approach enables a faster/slower adaptation when efficacy and safety endpoints are temporally aligned/misaligned, respectively.
-Keywords Project Optimus  $\cdot$ Patient-driven  $\cdot$ Quality Of Life  $\cdot$ Efficacy  $\cdot$ Adverse events  $\cdot$ Dose reduction  $\cdot$ Drug discontinuation  $\cdot$ Pharmacokinetic  $\cdot$ Pharmacodynamic  $\cdot$ Exposure–response
+(June 17, 2025)
+The majority of response-adaptive randomisation (RAR) designs in the literature rely on efficacy data to guide dynamic patient allocation.
+However, their applicability becomes limited in settings where efficacy outcomes, such as overall survival, are observed with a random delay.
+To address this limitation, we introduce SAFER (Safety-Aware Flexible Elastic Randomisation), a novel RAR design that leverages early-emerging safety data to inform treatment allocation decisions, particularly in oncology trials.
+The design is broadly applicable to contexts where prioritising the arm with a superior safety profile is desirable.
+This is especially relevant in non-inferiority trials, which aim to demonstrate that an experimental treatment is not inferior to the standard of care, while potentially offering improved safety or tolerability.
+In such trials, a well-established and unavoidable trade-off arises: maintaining statistical efficiency for the primary non-inferiority hypothesis while integrating safety-driven adaptations through RAR.
+The SAFER design addresses this trade-off by dynamically adjusting the allocation proportion based on the observed association between safety and efficacy endpoints.
+We illustrate the performance of SAFER through an extensive simulation study inspired by the CAPP-IT Phase III oncology trial.
+Results show that SAFER preserves statistical power when the safety-efficacy association is weak or absent and yields power gains when a strong positive association exists.
+Furthermore, SAFER reduces the adverse event rate during follow-up and offers flexible adaptation speed depending on the temporal alignment of the endpoints, faster when aligned and slower when misaligned.
+Keywords FDA Project Optimus, Efficacy, Toxicity, Adverse events, Endpoints association, Patient-benefit
 1 Introduction
@@
 Traditionally, early-phase clinical trials primarily prioritise safety, often focusing on major adverse events, with efficacy being largely exploratory.
-As development progresses, later phases are driven by efficacy decisions, and while comprehensive safety data are collected, this information is not typically formally integrated into design decisions, for example through an adaptive design.
-This leads to a critical challenge: despite the goal of improving the overall benefit-risk profile and enabling patients to stay on treatment longer, late-phase designs often do not formally use accumulating safety data to adapt trial conduct to be better aligned with those aims.
-The aforementioned issue becomes particularly relevant in the context of the modern oncology treatment landscape, which has been transformed by the advent of targeted therapies, monoclonal antibodies, and immune checkpoint inhibitors.
-These novel agents, designed to modulate specific molecular pathways, frequently demonstrate plateauing efficacy at doses well below the maximum tolerated dose (MTD) [2, 3].
-Moreover, their administration typically spans prolonged treatment cycles and is often accompanied by chronic, low-grade adverse events (AEs), including fatigue, gastrointestinal disturbances, and dermatologic reactions.
+As development progresses, later phases are primarily driven by efficacy decisions.
+Although comprehensive safety data are collected, this information is typically not formally integrated into design decisions, despite the goal of improving the overall benefit-risk profile and enabling patients to remain on treatment longer.
+This becomes particularly relevant in the context of the modern oncology treatment landscape, which has been transformed by the advent of targeted therapies, monoclonal antibodies, and immune checkpoint inhibitors.
+In fact, their administration typically spans prolonged treatment cycles and is often accompanied by chronic, low-grade adverse events (AEs), including fatigue, gastrointestinal disturbances, and dermatologic reactions.
 These AEs, although not classified as dose-limiting in the short windows of phase I trials, may significantly impair patient adherence, quality of life, and long-term tolerability in the later phases [2, 3].
+Moreover, novel agents, designed to modulate specific molecular pathways, often exhibit plateauing efficacy at doses well below the maximum tolerated dose (MTD) [4, 5], underscoring the need for a revised perspective on the interplay between safety and efficacy outcomes.
 A pertinent example to illustrate this is sotorasib (LUMAKRAS), a KRAS G12C inhibitor approved for non-small cell lung cancer [6].
@@
 Subsequent analyses showed that a 240 mg dose offered similar efficacy with a better safety profile, leading the FDA to mandate a post-marketing dose optimization study.
-The above rationale motivates our interest in the following crucial question: how can a late phase adaptive design help identify alternative dosing strategies—such as lower or fractionated doses, different schedules or regiments—deliver, with comparable or superior efficacy with fewer side effects?
+The above rationale motivates our interest in the following crucial question: how can a late phase adaptive design help identify alternative dosing strategies such as lower or fractionated doses, different schedules or regiments deliver, with comparable or superior efficacy and fewer side effects?
 This is particularly relevant to address the potential impact on phase II and III studies from dosing decisions from early trials often adopted on immediate and severe safety data.
 This is especially the case for agents with delayed efficacy signals and cumulative toxicity profiles that influence long-term adherence and outcomes.
-In light of the growing emphasis on toxicity management in oncology, we anticipate an increasing number of studies, both in early and late phase of development, will focus on evaluating alternative dosing regimens with improved safety profiles [7].
 Our work aims to address the research question presented above by leveraging Response-Adaptive Randomisation (RAR).
@@
 This adaptability is particularly advantageous when outcomes are available shortly after treatment initiation, allowing real-time modification of allocation proportions.
-However, in trials where primary efficacy outcomes are delayed – such as overall survival in oncology – RAR becomes more challenging due to the unavailability of timely, informative data during enrolment.
+However, in trials where primary efficacy outcomes are delayed, such as overall survival in oncology, RAR becomes more challenging due to the unavailability of timely, informative data during enrolment.
 Several strategies have been proposed to address this, including the use of early surrogate endpoints that correlate with long-term outcomes [9, 10, 11].
-On the contrary, this paper introduces a RAR design that, alternatively, uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
-However, a critical component of any safety-driven RAR is the inclusion on the possible interplay between safety and efficacy endpoints.
-While some drugs can maintain or improve efficacy after dose reductions, others could show reduced therapeutic benefit.
-From a statistical perspective, RAR designs may involve a trade-off with statistical power.
-In fact, while equal allocation does not necessarily maximise power in superiority trials [13], a balanced 1:1 allocation ratio is generally optimal in non-inferiority studies, where the alternative hypothesis assumes equal means (and variances) in the control and experimental arms.
+This paper introduces a RAR design that, alternatively, uses early-emerging safety data to inform treatment allocation decisions in oncology trials.
+However, a critical component of a safety-driven RAR design is the consideration of the potential interplay between safety and efficacy endpoints.
+While some drugs may maintain or even improve efficacy after dose reductions, others could exhibit a diminished therapeutic benefit.
+From a statistical standpoint, RAR designs may introduce challenges in preserving statistical power, especially under certain conditions.
+Although equal allocation does not universally guarantee maximal power in superiority trials, as highlighted by recent work by Pin et al.
+(2024) [12], it remains an effective and commonly adopted strategy in many trial settings.
+On the contrary, in the context of non-inferiority trials with binary (Bernoulli) or time-to-event (exponential) outcomes, a balanced 1:1 allocation ratio is typically optimal due to the mathematical structure of the variance, which is a function of the mean.
+Under the alternative hypothesis in such trials, the non-inferiority assumption postulates that the average treatment effects across the intervention and control arms are clinically equivalent, within a narrowly defined margin of non-inferiority.
+This margin is usually small [13], and as a result, the outcome variances in both arms are expected to be similar.
+Given this, allocation strategies that aim to minimize the variance of the estimated treatment effect, such as Neyman allocation [14, 15], which assigns subjects in proportion to the standard deviation of the outcome in each arm, tend to approximate or even coincide with balanced allocation in these settings.
+Thus, attempts to skew allocation dynamically based on accumulating data, as done in RAR designs, may deviate from this optimal variance-reducing strategy.
+Consequently, an unavoidable and well-established trade-off arises for such designs: to balance the goals of preserving inferential efficiency for the primary non-inferiority outcome while incorporating safety considerations into the randomisation process through RAR.
 To this regards, Baldi Antognini et al.
@@
 Although their work is aligned with our rationale, it is limited to single-outcome efficacy trials since it does not capture the complexity of a joint safety-efficacy decisions.
-Therefore, we propose an innovative Safety-Aware Flexible Elastic Randomisation (SAFER) design.
-SAFER dynamically adjusts patient allocation proportions based on accumulating safety endpoint data, aiming to strike a balance between patient safety, maximization of clinical benefit, and the exploitation of the association between efficacy and safety outcomes.
-This adaptive strategy is particularly suited for late-phase clinical trials, where it enhances decision-making while preserving statistical efficiency, especially with respect to study power.
+Therefore, we propose an innovative Safety-Aware Flexible Elastic Randomization (SAFER) design.
+SAFER dynamically adjusts patient allocation proportions based on accumulating safety endpoint data, with the aim of leveraging the association between efficacy and safety outcomes to balance clinical benefit and statistical validity.
+This adaptive strategy is particularly suitable for late-phase clinical trials, as it enhances patients’ treatment experience while preserving statistical power.
 The remainder of this paper is structured as follows.
-In Section 2, we define the key components of the SAFER design: (i) the efficacy endpoint, the associated power function, and sample size calculation; (ii) the definition of the safety endpoint; (iii) the target allocation proportion and the RAR procedure; and (iv) the integration of efficacy and safety data for the estimation of the allocation proportion.
+Section 2 introduces the key components of the SAFER design: (i) the efficacy endpoint, including the associated power function and sample size calculation; (ii) the definition of the safety endpoint; (iii) the target allocation proportion; and (iv) the integration of efficacy and safety data to guide the targeted allocation proportion.
 Section 3 introduces the CAPP-IT study as a motivating example and presents its re-design in a non-inferiority framework using the SAFER approach.
-Specifically, we outline the rationale and report the results from six distinct simulation scenarios, aimed at evaluating the performance of the proposed design under varying conditions.
+Specifically, we outline the rationale and report the results from several distinct simulation scenarios, aimed at evaluating the performance of the proposed design under varying conditions.
 Section 4 discusses the simulation findings, explores broader implications of the SAFER design, and concludes with a summary of key contributions and directions for future research.
-2 Components of the SAFER design
-2.1 Efficacy Endpoint (Power and Sample Size calculation)
+2 SAFER design components
+2.1 Efficacy Endpoint
 Our proposed design is specifically tailored to accommodate for the typical primary endpoint used in Phase II/III oncology trials.
@@
 Consequently, both the sample size and the primary analysis of our design will be expressed throughout this paper in terms of this designated endpoint.
-The sample size will be determined accordingly to ensure adequate statistical power for that main primary analysis.
+The sample size will be determined accordingly to the expected effect size to ensure adequate statistical power for that main primary analysis.
 We will now formally outline this rationale, including the underlying assumptions and notation.
-Let $j=E,C$ the experimental treatment and control arms, respectively.
+The exponential distribution was assumed for simulation purposes; the implications and limitations of this assumption are further discussed in the Discussion section.
+Let $j=E,C$ denote the experimental treatment and control arms, respectively.
 The patients’ primary efficacy endpoint (e) i.e., their survival time ($Y^{(e)}$ ) is assumed to be independent and identically distributed (i.i.d.) with probability density function (p.d.f) when treatment $j$ is assigned:
@@
 Therefore, $\theta_{j}^{(e)}$ represents the mean of the treatment effect, and $\theta_{j}^{(e)2}$ represents the variance of the treatment effect.
-We will focus our design on the case in which the study will be testing for the non-inferiority of E with respect to C, the non-inferiority hypothesis for the mean difference becomes [15, 16]
+We focus on the case in which the study is testing for the non-inferiority of the experimental with respect to the control arm.
+The null hypothesis for the mean difference can be expressed as [17, 18]
 $H_{0}:\theta_{C}^{(e)}-\theta_{E}^{(e)}\geq m\quad H_{1}:\theta_{C}^{(e)}-\theta_{E}^{(e)}<m,$ (2)
@@
 Based on this hypothesis, and under the assumptions that the means are strictly positive, and the standard regularity conditions hold (to ensure the asymptotic normality of the Maximum Likelihood Estimations), we can construct a Wald statistic as follows [16]:
-$W=\frac{\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)}-m}{\sqrt{Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})}}\sim\mathcal{N}(0,1),$ (3)
+$W=\frac{\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)}-m}{\sqrt{Var(\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)})}}\sim\mathcal{N}(0,1),$ (3)
 where the variance is estimated by
-$Var(\theta_{C}^{(e)}-\theta_{E}^{(e)})=\frac{1}{\sqrt{E}}\sqrt{\hat{\theta}_{C}^{(e)2}+\hat{\theta}_{E}^{(e)2}}$ (4)
+$\sqrt{Var(\hat{\theta}_{C}^{(e)}-\hat{\theta}_{E}^{(e)})}=\sqrt{N^{-1}\cdot(\hat{\theta}_{C}^{2}+\hat{\theta}_{E}^{2})},$ (4)
+with $N$ equal to the sample size.
+Although both the semi-parametric Cox model and the parametric exponential survival models are consistent estimators under exponentiality, in practice, the semi-parametric model often converges more quickly and accurately in simulations [19, 20].
+This is primarily due to its reliance on partial likelihood, fewer distributional assumptions, and better numerical conditioning in the presence of non-informative censoring and finite samples.
+Therefore, we estimate the Wald statistic through the Cox model [21].
 We also incorporate an early stopping feature into our designs, recognizing that this feature, when combined with RAR, can offer advantages over RAR alone [22].
 For simplicity, we have focused on only two analyses: an interim analysis (k=1) and the final analysis (k=2).
-Therefore, we will provide $W_{k=1}$ and $W_{k=2}$, estimated based on the first group available at the interim analysis and the complete sample at the end of follow-up, respectively.
-We estimate the treatment effect through the Cox model [18] such that $W$ is approximated by $(\hat{\beta}_{E}-m)/SE(\hat{\beta}_{E})$ where $\hat{\beta}_{E}$ is the estimated difference between arms, i.e., $(\hat{\theta}_{C}^{(p)}-\hat{\theta}_{E}^{(p)})$ and $SE(\hat{\beta}_{E})$ is the square root of $Var(\hat{\beta}_{E})$.
-Although both the Cox model and exponential survival models are consistent estimators under exponentiality, in practice, the Cox model often converges more quickly and accurately in simulations.
-This is primarily due to its reliance on partial likelihood, fewer distributional assumptions, and better numerical conditioning in the presence of non-informative censoring and finite samples [19, 20].
-Moreover, to estimate the minimum number of observed progression events $P$ needed for the study, we construct a $H_{0}$ rejection region as $R=\left\{-\infty,z_{1-\beta}\right\}$, with
+Therefore, we will provide test statistics $W_{k=1}$ and $W_{k=2}$, estimated based on the first group available at the interim analysis and the complete sample at the end of follow-up, respectively.
+To estimate the minimum number of observed progression events $P$ needed for the study, we construct a $H_{0}$ rejection region as $R=\left\{-\infty,z_{1-\beta}\right\}$, with
 $z_{1-\beta}=\sqrt{P\cdot(\log(HR_{H_{1}})-m)^{2}\pi_{E}\cdot(1-\pi_{E})}-z_{1-\alpha},$ (5)
-where $\alpha$ is the significance level, $\beta$ is the Type II Error rate, $HR_{H_{1}}$ is the hazard ratio ($\lambda_{E}/\lambda_{C}$) under the alternative hypothesis, and $\pi_{E}$ is the realized allocation proportion to the experimental arm (E).
-We will let $\pi_{E}$ vary from 0.5 to 0.8.
-Therefore, for each $\pi_{E}$ and associated rejection region, we can calculate the minimum number of events $P$ needed to test the non-inferiority on the primary endpoint between treatments, based on the following power function definition [21].
-[…]
-[…]
-After the calculation of the minimum number of events $P$ needed to achieve a study power of at least 0.8, we adjust the total number of events needed due to the application of the Lan-DeMets O’Brien-Fleming $\alpha$-spending function [22] to account for $k=2$ analyses.
-Finally, due to the survival primary endpoint, the overall sample size ($N$) is estimated by dividing the total number of events needed ($P$) by the probability to observe the event during the follow-up, given the event rate expected in the control and experimental arms, under the alternative hypothesis.
+where $\alpha$ is the significance level, $\beta$ is the type-II Error rate, $HR_{H_{1}}$ is the hazard ratio ($\lambda_{E}/\lambda_{C}$) under the alternative hypothesis, and $\pi_{E}$ is the allocation proportion to the experimental arm (E).
+In Section 3 we vary $\pi_{E}$ from 0.5 to 0.8.
+Therefore, given the rejection region, we can calculate the minimum number of events $P$ needed to test the non-inferiority on the primary endpoint between treatments, based on the following power function definition [23].
+$Power=P_{H_{1}}(Z_{1}\in\mathcal{R}_{1})+P_{H_{1}}(Z_{1}\in\mathcal{C}_{1},Z_{2}\in\mathcal{R}_{2})$ (6)
+In equation (6), $Z_{1}$ and $Z_{2}$ are the statistic tests at the interim and final analyses.
+$P_{H_{1}}(Z_{1}\in\mathcal{R}_{1})$ is $power_{1}$ (interim analysis power) i.e., the probability at the interim, under the alternative hypothesis ($H_{1}$), to fall into $R_{1}$, the region of rejection of the null hypothesis ($H_{0}$).
+Moreover, $P_{H_{1}}(Z_{1}\in\mathcal{C}_{1},Z_{2}\in\mathcal{R}_{2})$ is $power_{2}$ (final analysis power) i.e., the probability, under $H_{1}$ of reject $H_{0}$ at the final analysis given that the study continued at interim.
+Therefore, $power_{2}$ is given by a bivariate integral where $R_{2}$ and $C_{1}$ ($1-R_{1}$) indicate the rejection regions of $H_{0}$ and continuation at the final and interim analyses, respectively.
+After the calculation of the minimum number of events $P$ to reach an overall power of at least 0.8 (assuming $\alpha$ = 0.05, $\pi_{E}$ = 0.5, the $H_{1}$ and the $m$ as defined below in Section 3), we adjust the total number of events needed due to the application of the Lan-DeMets O’Brien-Fleming $\alpha$-spending function [24] to account for $k=2$ analyses.
+Finally, due to the survival primary endpoint, the overall sample size ($N$) is estimated by dividing the total number of events needed ($P$) by the probability to observe at least one event during the follow-up, i.e., $Pr(\text{at least one event in time }t)=1-e^{-\lambda\cdot t}$ where lambda is the weighted (across arms) average event rate expected in the sample and t the maximum follow-up time.
 2.2 Safety Endpoint
 The safety endpoint in our design is the time to dose-reduction/drug discontinuation due to low drug tolerability or AEs.
-The patient’s safety endpoint (s) ($Y^{(s)}$), is assumed to be i.i.d., with p.d.f.
+The patients’ safety endpoint (s) ($Y^{(s)}$), is assumed to be i.i.d., with p.d.f.
 when treatment $j$ is assigned:
@@
 The safety model and its estimation will directly inform the RAR component of the design, as detailed in the next subsection.
-2.3 Allocation Proportion Target
+2.3 Allocation proportion
 For our proposed design, the Neyman allocation [14, 15, 7] is a reasonable choice of allocation proportion to target.
-This is because Neyman allocation under these assumptions will consistently assign the majority of patients to the treatment that patients are observed to discontinue later [14] (or tolerate for the longest).
-The Neyman allocation proportion (or the proportion to the treatment arm $\pi_{E}$) can be estimated based on estimates of the safety average time to dose reduction as indicated below:
+First, the Neyman allocation, maximizes the statistical power of a widely used Wald test to detect the true difference between arms; second, under the assumption of exponentially distributed survival endpoint, the Neyman proportion will consistently assign the majority of patients to the treatment that patients are observed to discontinue later [16] (or tolerate for the longest).
+Note that the latter is not always true for other endpoint distributions, such as the binary case [25], where the Neyman allocation can favour the worst treatment to minimize the variability of the Wald test.
+In our design, the Neyman target allocation proportion is estimated based on the estimated safety average time to dose reduction as indicated below:
 $\hat{\pi}_{E}=\frac{\hat{\theta}_{E}^{(s)}}{\hat{\theta}_{E}^{(s)}+\hat{\theta}_{C}^{(s)}}$ (8)
-In our simulation design, we allow for $u=1,…,U$ update times for $\hat{\pi}_{E}$, which occur independently of the $k$-analyses defined by the group-sequential design (where interim efficacy analyses are assumed to occur only once, at $k=1$).
-Specifically, the randomisation proportion $\hat{\pi}_{E}$ is updated sequentially:
-  • For patients enrolled before the first update time ($u=1$), $\hat{\pi}_{E}$ is set to 0.5. This is the burn-in stage of the RAR feature of this design [26].
+It is worth highlighting that we define $\pi_{E}$ as the true target allocation proportion, computed by plugging the true values of $\theta_{E}^{(s)}$ and $\theta_{C}^{(s)}$ into equation 8, in accordance with the simulation settings used for safety data generation.
+In contrast, $\hat{\pi}_{E}$ denotes the estimated target allocation proportion, obtained by substituting $\hat{\theta}_{E}^{(s)}$ and $\hat{\theta}_{C}^{(s)}$ i.e., the estimations of the respective parameters, into the same equation.
+Specifically, in our simulation design, we allow for $u=1,\ldots,U$ update times for $\hat{\pi}_{E}$, which occur independently of the $k$-analyses defined by the group-sequential design (where interim efficacy analyses are assumed to occur only once, at $k=1$).
+Therefore, the randomisation proportion $\hat{\pi}_{E}$ is updated sequentially:
+  • For patients enrolled before the first update time ($u=1$), $\hat{\pi}_{E}$ is set to 0.5. This is the burn-in stage of the RAR feature of this design [26, 27].
   • For patients enrolled between update time $u=1$ and $u=2$, $\hat{\pi}_{E}$ uses the estimate of $\hat{\theta}_{E}^{(s)}$ and $\hat{\theta}_{C}^{(s)}$ from the first update time.
@@
 The number of update times ($U$) and the corresponding calendar occasions should be specified in the study protocol.
-The updates may be equally spaced across the enrolment period or not, depending on the design objectives.
-However, in the present design, we did not investigate how varying the number and timing of updates affects the design characteristics.
-Instead, we fixed equally spaced updates every three months, starting from month 3 through month 45, assuming a total enrolment period of 48 months.
+The updates may be equally spaced across the enrolment period or not, depending on the design objectives and practical considerations.
+We fixed equally spaced updates every three months, starting from month 3 through month 45, assuming a total enrolment period of 48 months.
 Crucially, each $\hat{\pi}_{E}$ is estimated using all available data up to a specific update time.
-2.3.1 Elastic Adaptation of Allocation Proportion: Integration of Efficacy and Safety
-A final element involves a function designed to modulate the RAR adaptation based on safety, utilizing the estimated association between safety and efficacy endpoints.
-The underlying idea is that if there’s sufficient evidence of non-inferiority, we would favour allocating more patients to the better-tolerated arms.
-Conversely, if non-inferiority is not supported by the current evidence in the study, the adaptation would shift away from those arms.
-This is achieved by the inclusion in the design of an elastic function that dynamically adjusts the estimated allocation target proportion $\hat{\pi}_{E}$ based on the observed treatment effect on PFS, during the enrolment period.
-The SAFER function is defined in equation (9).
-$SAFER(\hat{\pi}_{E})=\begin{cases}0.5&\text{if }\hat{\Phi}\leq 0.5\\ 0.5+(\hat{\pi}_{E}-0.5)\cdot\left(1-\left(1-\frac{\hat{\Phi}-0.5}{0.5}\right)^{\eta}\right)&\text{if }0.5<\hat{\Phi}<1\\ \hat{\pi}_{E}&\text{if }\hat{\Phi}=1\end{cases}$ (9)
-Note that the maximum possible adjusted allocation target to be used in the design is equal to the standard adaptive allocation proportion $\hat{\pi}_{E}$ and the minimum adjusted allocation proportion is set to 0.5, representing equal allocation.
-The intermediate cases are introduced through a tuning parameter $\hat{\Phi}$ that quantifies the association between the safety and efficacy endpoints.
-Specifically, $\hat{\Phi}$ is estimated using the cumulative density function (CDF) of the absolute z-value, derived from a survival regression model, which corresponds to the estimated treatment effect on PFS.
-We only consider $\hat{\Phi}\geq 0.5$, effectively excluding the possibility of a negative association between treatment and PFS.
-Moreover, the parameter $\eta\geq 1$ is the elastic parameter which controls the curvature of the function: $\eta=1$ corresponds to a linear function, while $\eta>1$ results in a concave shape, as illustrated in Figure 1.
+2.3.1 Efficacy and safety association
+To incorporate safety into RAR, the design includes a modulation function that dynamically combines emerging evidence on safety and efficacy.
+When the data indicate that the experimental treatment is better tolerated and that this improved safety profile translates into a favourable efficacy outcome, the algorithm increases the probability of assigning future patients to the experimental arm.
+If the data do not confirm an efficacy benefit despite the safety advantage, randomisation resets to an equal (1:1) allocation between the two arms.
+This adaptive mechanism is governed by a parameter $\hat{\Phi}$, which quantifies the strength of the evidence supporting the impact of improved safety on efficacy.
+It is defined as the cumulative distribution function (CDF) of the $z$-value associated with the estimated treatment effect on PFS, obtained from the survival model.
+In fact, a simple correlation metric between safety and efficacy is not appropriate in this context, as it fails to account for the specific characteristics of the survival endpoint.
+In particular, both endpoints are subject to truncation: the time-to-event safety endpoint (e.g., time to dose reduction or drug discontinuation) and the PFS endpoint are both affected by administrative censoring.
+Moreover, the safety endpoint is inherently truncated by the occurrence of PFS events, since once progression occurs, the follow-up ends.
+As a result, the two endpoints are structurally correlated, and conventional correlation coefficients may yield misleading estimates of their association.
+Specifically, let:
+  • $\hat{\theta}_{\text{PFS}}$ denote the estimated treatment effect (e.g., log hazard ratio) on PFS;
+  • $\text{SE}(\hat{\theta}_{\text{PFS}})$ its standard error;
+  • $Z=\dfrac{\hat{\theta}_{\text{PFS}}}{\text{SE}(\hat{\theta}_{\text{PFS}})}$ the corresponding $z$-statistic;
+  • $\Phi(\cdot)$ the standard normal cumulative distribution function.
+Since we expect a hazard ratio lower than 1 in case of treatment effect, we take the absolute value of the z-statistic. Therefore, the adaptation weight is defined as:
+$\hat{\Phi}=\Phi\left(|Z|\right)=\Phi\left(\left|\dfrac{\hat{\theta}_{\text{PFS}}}{\text{SE}(\hat{\theta}_{\text{PFS}})}\right|\right)$ (9)
+Specifically, this parameter $\hat{\Phi}$ is incorporated into the allocation function as follows:
+  • When the experimental arm is better tolerated (i.e., the estimated Neyman allocation is greater than 0.5) and $\hat{\Phi}$ exceeds a predefined threshold, the allocation probability for the experimental arm is increased.
+  • When $\hat{\Phi}$ is below the threshold, suggesting insufficient evidence, allocation probabilities revert toward equal randomisation.
+Since the $\hat{\Phi}$ ranges from 0 to 1, the default threshold is set at 0.5 (corresponding to a z-statistic equal to zero).
+However, sensitivity analyses should be conducted to assess the effect of alternative choices.
+Therefore, the SAFER function is defined as:
+$\text{SAFER}(\hat{\pi}_{E})=\begin{cases}0.5&\text{if }\hat{\Phi}\leq 0.5\\ 0.5+(\hat{\pi}_{E}-0.5)\cdot\left(1-\left(1-\dfrac{\hat{\Phi}-0.5}{0.5}\right)^{\eta}\right)&\text{if }0.5<\hat{\Phi}<1\\ \hat{\pi}_{E}&\text{if }\hat{\Phi}=1\end{cases}$ (10)
+Note that the maximum allocation proportion used in the SAFER design corresponds to the estimated Neyman allocation $\hat{\pi}_{E}$ when $\hat{\Phi}=1$, while the minimum is fixed at 0.5, representing complete randomisation, when $\hat{\Phi}\leq 0.5$.
+Sensitivity analyses should also explore scenarios with upper bounds below 1 (e.g., 0.9) to assess robustness.
+Moreover, the parameter $\eta$ is the elastic parameter which controls the curvature of the function: $\eta=1$ corresponds to a linear function, while $\eta>1$ results in a concave shape, as illustrated in Figure 1.
 Although convex shapes ($\eta<1$) are theoretically allowed, we did not explore the values of $\eta<1$, as the linear function is already considered a conservative choice to adjust the allocation proportion in cases of small $\hat{\Phi}$ values.
 Figure 1: Target function of the SAFER design for two target allocations, $\hat{\pi}_{E}=0.7$ and $\hat{\pi}_{E}=0.9$, each evaluated under three shape parameters $\eta=1,2,5$.
-The SAFER function increases monotonically in $\Phi_{z}$ and reflects more aggressive allocation as both $\hat{\pi}_{E}$ and $\eta$ increase.
-3 Performance Evaluation Through Simulation: redesigning a non-inferiority study
+The SAFER function increases monotonically in $\hat{\Phi}$ and reflects more aggressive allocation as both $\hat{\pi}_{E}$ and $\eta$ increase.
+3 SAFER design evaluation
 This section details the extensive simulation studies conducted to systematically assess the performance of the proposed SAFER design methodology.
-These simulations aim to evaluate key operating characteristics including statistical power and allocation patterns under various scenarios.
+These simulations aim to evaluate key operating characteristics including statistical power and the estimated allocation proportion under various scenarios.
 The findings presented herein are based on a specific case study, which we redesign using SAFER and help us provide a comprehensive understanding of the SAFER’s design features and its potential advantages in different trial settings.
@@
 Moreover, we assume a fixed total duration of the trial equal to 60 months, considering a uniformly distributed enrolment of 48 months and 12 months of follow-up.
-Assuming a monthly event rate ($\lambda_{C}$) of 0.069, which corresponds to a median PFS of approximately 10 months, and a $HR_{H0}$ of 1.25 ($\lambda_{E}/\lambda_{C}$), the median PFS in the experimental arm must exceed 8 months to demonstrate non-inferiority with respect to the control arm.
-3.2 Modelling efficacy and safety endpoints association
-To thoroughly assess SAFER’s performance, understanding the interplay between primary efficacy and safety/tolerability endpoints is essential.
-This section details how this relationship is incorporated into our framework, specifically addressing scenarios where these endpoints are not assumed to be independent.
+Assuming a monthly event rate in the control arm ($\lambda_{C}$) of 0.069, which corresponds to a median PFS of approximately 10 months, and a $HR_{H0}$ of 1.25 ($\lambda_{E}/\lambda_{C}$), the median PFS in the experimental arm must exceed 8 months to demonstrate non-inferiority with respect to the control arm.
+3.2 Modelling the association between safety and efficacy endpoints
 […]
@@
 3.3 Metrics to assess performance
-Through simulations, we evaluated six distinct scenarios to quantify the trial’s performance under various conditions (described below), providing insights into its statistical properties and broader implications.
-For each scenario, we conducted 10,000 iterations, ensuring a Monte Carlo error of less than 0.5% for both power and type I error rate [28].
-We report the overall study power, calculated as the proportion of simulated trials that rejected $H_{0}$ under the alternative hypothesis ($H_{1}$), aggregated across all simulation studies.
-Moreover, the estimated allocation proportion to the experimental arm is defined as the average proportion of patients assigned to treatment $E$ at the end of each simulated trial.
-To assess the performance of the design and compare the SAFER approach with the fixed allocation proportion approach, we provide benchmark results—specifically, power and allocation proportions—obtained from a fixed randomisation design under various allocation proportions (from 0.5 to 0.8).
-These benchmarks were derived using a theoretical (non-simulation-based) approach, based on the normal approximation of the outcome as defined in equations (5) and (6), and are presented in Figure 3.
-The power loss is attributed to changes in the allocation proportion from 0.5 to 0.8.
-The overall power is maximized by $\pi_{E}=0.5$ since the $\lambda_{C}$ and $\lambda_{E}$ are assumed equal, and decreases to approximately 0.79 when $\pi_{E}=0.6$, 0.74 when $\pi_{E}=0.7$, and 0.63 when $\pi_{E}=0.8$.
-Figure 3: Theoretical loss of power, using normal approximation, for the primary endpoint analysis corresponding to diverse Information Fractions of 50% and allocation proportions $\pi_{E}$, from 0.5 to 0.8.
-Furthermore, to empirically verify the validity of the standard regularity conditions, we conducted a simulation using exponential distributions for survival outcomes, varying the information fraction (IF) at interim from 0.2 to 0.5; the results are reported in Table 1.
-The observed loss of power at increasing allocation proportions closely aligns with the theoretical results in Figure 3, suggesting that for further comparisons, the normal approximation can be reliably used in this setting.
-Interestingly, it seems that the IF does not modify the observed loss of power.
-Table 1: Power using exponential distributed outcomes with fixed allocation proportion and varying information fraction at interim analysis.
+We define a safety-driven RAR design as one in which the allocation proportion is updated solely based on the estimated Neyman allocation proportion.
+In contrast, we refer to the SAFER design when the allocation updates incorporate the SAFER targeting function, which accounts for the observed association between efficacy and safety.
+To compare these adaptive designs with a fixed allocation approach, we provide benchmark results focusing on statistical power.
+Specifically, we assess power using both a theoretical (non-simulation-based) approach as defined in equations (5) and (6) (see Figure A1) and a simulation-based approach, which employs exponentially distributed survival times as outcomes.
+The latter allows us to empirically assess the validity of the asymptotic assumptions underlying the theoretical approximation (see Table A1).
+Specifically, in Figure A1 the overall power is maximized by $\pi_{E}=0.5$ and decreases to approximately 0.79 when $\pi_{E}=0.6$, 0.74 when $\pi_{E}=0.7$, and 0.63 when $\pi_{E}=0.8$.
+In Table A1, the observed loss of power at increasing allocation proportions closely aligns with the theoretical results in Figure A1, suggesting that for further comparisons, the normal approximation can be reliably used in this setting.
+Interestingly, it seems that the Information Fraction (IF) at interim does not modify the observed loss of power.
+Moreover, through simulations, we evaluate several distinct scenarios to assess the design under various conditions (described below), providing insights into its statistical properties and broader implications.
+As metrics, we report the overall study power and type-I-error rate, calculated as the proportion of simulated trials that rejected $H_{0}$ under $H_{1}$ and $H_{0}$, respectively, the $power_{1}$, calculated as the proportion of simulated trials that rejected $H_{0}$ under $H_{1}$ at interim analysis (probability of early efficacy stop), the estimated allocation proportion to the experimental arm, defined as the average proportion of patients assigned to treatment $E$ at the end of each simulated trial, and the AE rate per arm calculated as the total number of AEs divided by the total person-years, i.e., the number of patients multiplied by the follow-up duration in years, to yield a rate per patient-year.
+For each scenario, we conducted 10,000 iterations, ensuring a Monte Carlo error of less than 0.5% for both power and type-I error rate control [29].
+3.4 Simulating scenarios
+In this section, we will describe in detail the distinct scenarios used for the simulation study.
+Scenario 0: Safety-driven RAR design (without a SAFER approach) under independence between the efficacy and safety endpoints.
+This scenario assesses the impact of a safety-driven RAR design—without implementing the SAFER approach—on power and type-I error rate control, under the assumption of independence between the efficacy and safety endpoints.
+The median time to dose reduction or drug discontinuation is simulated at 1.5 months in the control group, while in the experimental group, it is varied across values of 1.5, 2.25, 3.5, and 6 months corresponding to $\pi_{E}$ equal to 0.5, 0.60, 0.70, and 0.80, respectively.
+Moreover, we set $U=15$, resulting in updating $\hat{\pi}_{E}$ every three months until month 45, following an initial burn-in period during which a 1:1 randomization is applied for the first three months (about 60 patients).
+Scenario 1: Complete randomization design under diverse level of association between the efficacy and safety endpoints.
+This scenario is designed to assess the impact of diverse level of association between the efficacy and safety endpoints on statistical power and AEs rate.
+Specifically, we simulate increasing levels of association between the endpoints, with $E[\gamma_{1}]$ set to 0.001, 0.005, 0.01, 0.03, and 0.05, corresponding to very weak, weak, moderate, strong, and very strong associations, respectively.
+Moreover, we set the median AEs time in each arm corresponding to $\pi_{E}$ from 0.5 (no difference in safety between arms) to 0.8.
+To evaluate the effect of association between endpoint at the net of RAR, we simulate under a 1:1 allocation ratio (fair coin toss).
+Scenario 2: Safety-driven RAR design (without a SAFER approach) under diverse level of association between the efficacy and safety endpoints.
+This scenario is designed as a counterpart to scenario 1 and aims to assess the impact on power, allocation proportion, and AEs rate of the safety-driven RAR design.
+Therefore, at each time point $u=1,\ldots,U$, we compute $\hat{\pi}_{E}=\hat{\theta}_{E}^{(s)}/(\hat{\theta}_{E}^{(s)}+\hat{\theta}_{C}^{(s)})$ and use it as the allocation probability to arm $E$ for the randomization of patients in the subsequent three months.
+Under this scenario, we expect power to decrease in the presence of weak association between endpoints, and to increase as association strengthens, potentially due to the enhanced effectiveness of the experimental treatment on the efficacy endpoint driven by prolonged well-tolerated drug exposure.
+Scenario 3: SAFER design: impact on power under diverse level of association between the efficacy and safety endpoints, varying the $\eta$ values.
+This scenario is designed to evaluate the impact on power, allocation proportion, and AEs rate of the SAFER design, by varying the elastic parameter $\eta$ values.
+Specifically, scenarios 3a and 3b report the results of simulations conducted with $\eta=1$ and $\eta=5$, respectively.
+These values correspond to the absence of elasticity and a high degree of elasticity in the target SAFER function, as reported in Figure 1.
+Scenario 4: Impact of differential timing in the observation of endpoint data on $SAFER(\hat{\pi}_{E})$.
+This scenario aims to assess the speed at which $SAFER(\hat{\pi}_{E})$ reaches the target level, when the median times of the safety and efficacy endpoints are either close or far apart i.e., delayed efficacy outcome [10].
+For instance, we expect a faster pattern when the median PFS is close to the median time to dose reduction or drug discontinuation because there will be enough PFS events for the $\hat{\Phi}$ parameter estimation, early in the recruitment period.
+On the other hand, when evidence of the positive association between safety and efficacy emerges later in the recruitment, we expect the SAFER function to down-weight the allocation proportion for a longer period until sufficient evidence on PFS becomes available.
+Under this scenario, the median PFS ranges from 3 to 24 months.
+Scenario 5: Impact of informative drop-out (related to the safety endpoint) on $SAFER(\hat{\pi}_{E})$ and power.
+This scenario aims to assess the impact of informative drop-out on the operating characteristics of the SAFER design.
+We define informative drop-out as a situation in which a patient is observed until experiencing a side effect and subsequently drops out of the study.
+In practice, this means that the patient is followed until dose reduction or drug discontinuation occurs, after which neither the time to PFS nor the occurrence of the progression event is observed.
+We simulated the average allocation proportion and power under varying informative drop-out rates, ranging from 5% (very low) to 25% (very high), with drop-outs uniformly distributed over the accrual period.
+To address the potential bias introduced by informative drop-out, we considered the strategies for handling intercurrent events as defined in the Estimands framework [30].
+Specifically, we considered the composite strategy, where the time-to-PFS is defined as the last available follow-up time, with the occurrence of the progression = yes.
+Scenario 6: Impact of under-reported safety events on $SAFER(\hat{\pi}_{E})$ and power.
+This scenario aims to assess the impact of under-reported safety events on the operating characteristics of the SAFER design.
+Unlike the informative drop-out phenomenon—where patients are observed until an adverse event occurs and then drop out—in the case of under-reported safety events, it is not recorded at all and only the survival outcome (i.e., progression and the corresponding time-to-PFS) is available [31].
+3.5 Simulation Results
+The total sample size required to achieve at least a power value of 0.8 is 881 (about 18 patients enrolled per months).
+However, due to the interim analysis, the total number of events needed increases to 888 when IF = 0.5.
+Table 1 presents the results of the simulations under Scenario 0 (independence between endpoints).
+Both the estimated allocation proportion of patients assigned to the experimental arm $N_{E}/N$ and the power closely align with the target allocation $\pi_{E}$ and the theoretical power reported in Table A1, respectively.
+Minor discrepancies can be attributed to the burn-in period, as Table A1 reports results based on unequal allocation without any burn-in, whereas Scenario 0 includes this initial adjustment phase.
+The type-I error rate remains close to the nominal value of 0.05, with minor increases.
+However, these increases do not show a consistent pattern and are likely due to simulation variability.
+Therefore, neither the safety-driven RAR design nor the information fraction (IF) appears to negatively affect type-I error control under the null hypothesis.
+Table 1: Estimated allocation proportion, power, and type-I error rate summary: Scenario 0
+Allocation Power/Type-I error rate Information Fraction
+$\pi_{E}$ $N_{E}/N$ 0.2 0.3 0.4 0.5
+0.5 0.5 Power 0.80 0.80 0.80 0.80
+0.5 Type-I error rate 0.048 0.050 0.048 0.053
+0.6 0.59 Power 0.78 0.79 0.78 0.78
+0.59 Type-I error rate 0.047 0.050 0.055 0.053
+0.7 0.69 Power 0.74 0.74 0.73 0.73
+0.69 Type-I error rate 0.049 0.051 0.051 0.053
+0.8 0.78 Power 0.66 0.66 0.65 0.66
+0.78 Type-I error rate 0.050 0.055 0.056 0.051
+$\pi_{E}$ = True target allocation value by simulation setting; $N_{E}/N$ = Proportion of patients assigned to the experimental arm using a safety-driven RAR design.
+Table 2 presents the results of the simulations under Scenarios 1, 2, 3a, and 3b.
+Overall, it shows how the SAFER method effectively balances the benefits of a safety-driven RAR design with the need to maintain statistical validity.
+Scenario 1 is simulated using a complete randomization approach.
+Therefore, the observed increase in both $power_{1}$ and overall power can be attributed to the influence of the safety endpoint on the efficacy endpoint.
+Specifically, the longer the time to AEs in arm E, the longer the PFS tends to be in the same arm, which increases the probability of rejecting the null hypothesis.
+This trend is amplified by the strength of the association between the endpoints, with $power_{1}$ reaching 0.82 when $\pi_{E}=0.8$ and the association is very strong.
+Scenario 2 is simulated using a safety-driven RAR design, without the SAFER approach.
+As the estimated allocation proportion $N_{E}/N$ increases, both $power_{1}$ and the overall power decrease in the presence of very weak or weak associations between endpoints.
+Conversely, when the strength of association increases, a gain in power is observed.
+This improvement is driven by the meaningful positive effect on PFS resulting from drug exposure free from AEs, as described in Section 3.2.
+Under Scenario 3a, the SAFER design is implemented with the elastic parameter $\eta$ fixed at 1.
+In the presence of weak endpoint associations, the SAFER approach effectively shrinks the estimated allocation proportion towards 0.5, thereby restoring statistical power while still favouring the better-tolerated arm—though to a lesser extent.
+Conversely, when the association between endpoints is strong, the design allows higher allocation proportions in favour of the experimental arm without compromising statistical power.
+A similar pattern is observed in Scenario 3b, where the higher value of the elastic parameter $\eta=5$ allows for a more aggressive approach in reaching the target allocation proportion.
+This comes at the cost of a slight reduction in power compared to Scenario 3a.
+A crucial metric for evaluating the benefit of the proposed approach is the reduction in the AE rate per patient-year.
+Specifically, under the CR design (Scenario 1), the AE rate decreases as $\pi_{E}$ increases, from 1.28 when $\pi_{E}=0.5$ to approximately 1.05 when $\pi_{E}=0.8$.
+However, thanks to the safety-driven RAR design in Scenario 2 and the SAFER design in Scenarios 3a and 3b, the probability of being assigned to the better-tolerated arm increases.
+As a result, a larger reduction in the AE rate per patient-year is observed, with values dropping below 1 for high $N_{E}/N$ values and strong endpoint associations.
+Table 2: Estimated allocation proportion, power, and AEs rate summary: Scenarios 1, 2, 3a, and 3b (IF = 0.5)
 {tblr}
-cells = c, cell12 = c=4, vline2 = 1, vline2-5 = 2, hline1-3,10 = -, Allocation & Information Fraction
+column1 = c, column7 = c, column11 = c, column15 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4c, cell17 = c=4, cell111 = c=4, cell115 = c=4, cell23 = c, cell24 = c, cell25 = c, cell26 = c, cell31 = r=5, cell81 = r=5, cell131 = r=5, cell181 = r=5, hline1,3,8,13,18,23 = -, hline2 = 3-18, $\pi_{E}$ Endpoint
+association
+CR (1) Safety-driven RAR (2) SAFER (3a) SAFER (3b)
+$N_{E}/N$ $p_{1}$ $p$ AEr $N_{E}/N$ $p_{1}$ $p$ AEr $N_{E}/N$ $p_{1}$ $p$ AEr $N_{E}/N$ $p_{1}$ $p$ AEr
+0.5 Very Weak 0.50  0.21 0.80 1.28 0.50 0.21 0.80 1.28 0.50  0.21 0.80 1.28 0.50 0.21 0.80 1.28
+Weak 0.50 0.21 0.80 1.28 0.50 0.21 0.80 1.28 0.50 0.21 0.80 1.28 0.50 0.21 0.80 1.28
+Moderate 0.50 0.21 0.80 1.28 0.50 0.21 0.80 1.28 0.50 0.21 0.80 1.28 0.50 0.22 0.80 1.28
+Strong 0.50 0.21 0.80 1.27 0.50 0.20 0.78 1.27 0.50 0.21 0.79 1.27 0.50 0.21 0.79 1.27
+Very Strong 0.50 0.21 0.80 1.27 0.50 0.20 0.79 1.27 0.50 0.21 0.79 1.27 0.50 0.21 0.79 1.27
+0.6 Very Weak 0.50 0.21 0.80 1.23 0.59 0.22 0.79 1.22 0.52 0.22 0.81 1.23 0.54 0.22 0.80 1.23
+Weak 0.50 0.23 0.80 1.23 0.59 0.22 0.80 1.22 0.53 0.22 0.81 1.23 0.54 0.22 0.82 1.23
+Moderate 0.50 0.23 0.83 1.23 0.59 0.24 0.82) 1.22 0.53 0.25 0.84 1.23 0.54 0.25 0.84 1.23
+Strong 0.50 0.28 0.87 1.23 0.60 0.28 0.86 1.22 0.53 0.29 0.87 1.23 0.55 0.29 0.87 1.22
+Very Strong 0.50 0.33 0.91 1.23 0.60 0.31 0.90 1.22 0.53 0.33 0.91 1.22 0.55 0.32 0.91 1.22
+0.7 Very Weak 0.50 0.22 0.81 1.16 0.69 0.19 0.75 1.12 0.55 0.22 0.80 1.15 0.58 0.21 0.81 1.14
+Weak 0.50 0.25 0.84 1.16 0.69 0.21 0.78 1.12 0.55 0.24 0.84 1.15 0.59 0.24 0.84 1.14
+Moderate 0.50 0.27 0.87 1.16 0.69 0.24 0.81 1.11 0.56 0.27 0.87 1.15 0.59 0.27 0.87 1.14
+Strong 0.50 0.41 0.95 1.15 0.69 0.35 0.92 1.11 0.58 0.41 0.95 1.14 0.62 0.39 0.95 1.13
+Very Strong 0.50 0.55 0.98 1.15 0.69 0.49 0.96 1.11 0.60 0.54 0.98 1.13 0.64 0.53 0.99 1.12
+0.8 Very Weak 0.50 0.23 0.81 1.05 0.78 0.16 0.66 0.92 0.57 0.22 0.81 1.01 0.62 0.20 0.81 0.99
+Weak 0.50 0.28 0.87 1.05 0.78 0.19 0.73 0.92 0.59 0.27 0.87 1.01 0.64 0.25 0.86 0.98
+Moderate 0.50 0.34 0.92 1.04 0.78 0.24 0.79 0.91 0.60 0.33 0.92 1.00 0.66 0.31 0.91 0.97
+Strong 0.50 0.62 0.99 1.03 0.79 0.46 0.95 0.90 0.66 0.59 0.99 0.96 0.71 0.56 0.99 0.94
+Very Strong 0.50 0.82 1.00 1.03 0.79 0.66 0.99 0.89 0.70 0.80 1.00 0.93 0.74 0.76 1.00 0.91
+AEr = Adverse event rate; CR = Complete randomization; $p_{1}$ = Power at interim; $p$= Overall power; $\pi_{E}$ =True target allocation by simulation setting; $N_{E}/N$ = Proportion of patients assigned to the experimental arm.
+Figure 3 and Figure 4 illustrate the median allocation proportion, estimated using the SAFER design, over an accrual period of 48 months.
+Through scenario 4, we underscore the critical role of endpoint timing in the proposed design.
+When the safety and efficacy endpoints are temporally aligned (Panels A and B), the design detects associations more rapidly, enabling faster and more confident adjustments in allocation proportions.
+In contrast, greater temporal separation between endpoints (Panels C and D) delays the detection of such associations and slows the adaptation process.
+This finding highlights a key advantage of the SAFER design: the flexibility provided by the proposed function allows for better control over the degree of adaptation, particularly in settings with delayed efficacy outcomes, a common feature in survival studies.
+Unlike the safety-driven RAR approach, which update the allocation proportion based solely on the safety endpoint data, the SAFER design incorporates regularization that prevents premature and extreme imbalances in allocation.
+This is especially relevant in non-inferiority settings, where the experimental arm could ultimately prove inferior.
+By moderating allocation shifts until sufficient evidence accumulates on the efficacy endpoint, the design protects patients from being preferentially assigned to an arm that performs better on safety outcomes but may fail on efficacy assessment.
+Importantly, this guarantees that, under the null hypothesis, the allocation mechanism does not favour an inferior treatment, even if it appears more tolerable.
+Figure 3: $SAFER(\hat{\pi}_{E})$ vs.
+enrolment period, under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;).
+True target allocation by simulation setting=0.8; IF=0.5; $\eta$=5; Association between endpoints: very strong.
+Figure 4: $SAFER(\hat{\pi}_{E})$ vs.
+enrolment period, under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;).
+True target allocation by simulation setting=0.8; IF=0.5; $\eta$=5; Association between endpoints: weak.
+Further, the study assessed the impact of two common issues in oncology trials: Table 3 reports the estimated allocation proportion $SAFER(\hat{\pi}_{E})$ and the estimated power at increasing percentages of drop-out and under-reporting rates.
+A direct comparison with the values reported in Table 2 corresponding to scenario 3b and association "Very Strong" is possible since both the strength of the endpoint association and $\eta$ are fixed at the same levels.
+The adoption of a composite strategy to address informative drop-out (scenario 5) results in an increase in the estimated allocation proportion as the drop-out rate rises.
+Specifically, under this strategy, the time-to PFS is imputed as equal to the time-to-drug reduction for patients who discontinue prematurely.
+Consequently, the allocation proportion $N_{E}/N$ gradually converges toward the Neyman allocation proportion estimation $\hat{\pi}_{E}$.
+This convergence occurs because the observed differences in PFS between treatment arms become increasingly similar to the differences in time-to-drug reduction, thereby diminishing the influence of the down-weighting effect introduced by the SAFER approach.
+In parallel, statistical power also increases.
+This is attributable to the growing number of events contributing to the composite endpoint (safety plus efficacy) as the drop-out rate rises.
+Additionally, by design, the treatment arms show a substantial difference on the safety endpoint, which enlarge the overall effect size captured by the composite measure.
+As a result, the likelihood of rejecting the null hypothesis increases.
+While a higher drop-out rate introduces bias into the estimation process, under the alternative hypothesis, this bias may not necessarily be detrimental to patient outcomes.
+However, a problematic scenario may arise if the experimental arm demonstrates better tolerability but worse efficacy.
+In such cases, the composite strategy may fail to protect the efficacy endpoint analysis.
+Specifically, imputing the time-to-PFS as equal to the time of dose reduction or discontinuation may erroneously lead to the rejection of the null hypothesis, thereby substantially increasing the risk of type-I error rate inflation.
+Importantly, this bias is not introduced by the RAR design itself, but rather by the nature of the composite strategy—particularly when drop-out is more likely in the presence of poor tolerability, and tolerability differs significantly between arms.
+Therefore, when adopting a composite strategy, the treatment effect on the efficacy endpoint should be interpreted as a combined measure of both efficacy and safety since it reflects the integrated impact of tolerability on clinical outcomes.
+Under this interpretation only, the use of a composite strategy is reasonable and justified.
+As an alternative, when a biologically plausible and well-specified association exists between the efficacy and safety endpoints, a hypothetical strategy may be used.
+Within this framework, the occurrence of the efficacy endpoint can be modelled conditionally on the observed safety outcome, thereby estimating what would have occurred in the absence of the intercurrent event [32].
+This approach can preserve the Estimand consistency and potentially improve inference efficiency, provided the assumptions underlying the association model are adequately met.
+On the other hand, if the efficacy data are available even for patients who drop out, a treatment policy strategy could be adopted.
+This would mitigate the risk of biased inference due to the informative censoring [30, 32].
+In general, further analyses and comprehensive simulation studies are necessary to thoroughly assess the effects of informative drop-out under varying scenarios and parameter settings and the best way to handle the bias in this setting.
+On the other hand, in scenario 6, as the under-report rate increases, the allocation proportion decreases.
+This result is expected, as the core of the proposed RAR method relies on detecting differences in the safety endpoint.
+The greater the under-reporting of this endpoint, the less evidence the algorithm has to adjust the allocation proportion in favour of the better-tolerated arm.
+Table 3: Estimated allocation proportion and power summary: Scenarios 5 and 6 (IF=0.5, $E[\gamma]=0.05$, $\eta=5$)
+{tblr}
+width = colspec = Q[144]Q[119]Q[227]Q[90]Q[267]Q[90], columnodd = c, cell11 = r=2, cell12 = r=2, cell13 = c=20.317, cell15 = c=20.357, cell31 = r=5, cell81 = r=5, cell131 = r=5, cell181 = r=5, hline1,23 = -0.08em, hline2 = 3-6, hline3 = 1-2,4,6, hline3 = 3,50.03em, hline8,13,18 = 1-3,50.03em, hline8,13,18 = 4,6, $\pi_{E}$ Rate Informative drop-out (5) AEs under-reporting (6)
+$N_{E}/N$ $p$ $N_{E}/N$ $p$
+0.5 Very Low 0.50 0.81 0.50 0.79
+Low 0.50 0.81 0.50 0.79
+Moderate 0.50 0.83 0.50 0.79
+High 0.50 0.85 0.50 0.79
+Very High 0.50 0.87 0.50 0.79
+0.6 Very Low 0.55 0.92 0.55 0.91
+Low 0.56 0.94 0.54 0.91
+Moderate 0.56 0.95 0.54 0.91
+High 0.56 0.96 0.54 0.90
+Very High 0.56 0.97 0.53 0.91
+0.7 Very Low 0.64 0.99 0.63 0.98
+Low 0.65 0.99 0.62 0.98
+Moderate 0.65 1.00 0.61 0.99
+High 0.66 1.00 0.60 0.99
+Very High 0.67 1.00 0.60 0.98
+0.8 Very Low 0.75 1.00 0.73 1.00
+Low 0.76 1.00 0.72  1.00
+Moderate 0.76 1.00 0.71 1.00
+High 0.77 1.00 0.70 1.00
+Very High 0.77 1.00 0.69 1.00
+AE = Adverse event; $p$ = Overall power; $\pi_{E}$ = True target allocation by simulation setting; $N_{E}/N$ = Proportion of patients assigned to the experimental arm using SAFER approach.
+4 Discussion
+This paper introduces a safety-driven RAR design with the aim to enhance patient-oriented research by preferentially allocating participants to treatment arms with a more favourable safety profile.
+This approach is particularly relevant in oncology and may support the design of new trials to assess the long-term efficacy and tolerability of novel agents that are typically administered over prolonged treatment cycles and are frequently associated with chronic and cumulative toxicities—major contributors to treatment discontinuation and reduced efficacy.
+Moreover, a key feature of our design is the explicit evaluation of the interplay between efficacy and safety endpoints, an aspect increasingly emphasized by regulatory authorities and highlighted by the FDA through initiatives such as Project Optimus.
+Main Results
+Our results confirm the trade-off between safety-driven allocation design and statistical power for the efficacy endpoint analysis, in two-arm non-inferiority trials.
+Therefore, we propose a SAFER design to provide a practical and flexible framework for incorporating safety considerations into the randomisation process while maintaining statistical rigour, by dynamically adjusting the allocation proportion based on the observed association between safety and efficacy outcomes.
+Particularly, in the absence of such an association, the algorithm conservatively shrinks the allocation proportion toward 0.5; conversely, when a positive association between safety and efficacy is detected, whether strong or weak, the algorithm allows proportionally more aggressive adaptations.
+In doing so, the SAFER design preserves the statistical power of the primary analysis and adhering to the hierarchical endpoints structure and the non-inferiority objective: to identify a new treatment that, given clinically equivalent efficacy, can reduce the incidence of AEs, enhance the overall patient experience, and, in certain cases, even improve drug effectiveness through a more favourable therapeutic profile due to fewer side effects.
+Notably, under associated endpoints, statistical power is not only preserved but often increases with stronger association.
+This can be attributed to improved efficacy outcomes resulting from longer or more tolerable exposure to the experimental treatment.
+Considerations on design flexibility and generalizability
+The current design adopts a group-sequential framework with a single interim analysis and a final analysis.
+However, it is fully customizable, allowing for the specification of multiple interim analyses and the selection of various $\alpha$-spending functions.
+Additionally, a $\beta$-spending function can be incorporated to formally account for futility stopping, thereby providing greater flexibility in controlling both type-I and type-II error rates [33].
+Although the proposed design is developed within a frequentist framework, a Bayesian counterpart can be readily implemented.
+Specifically, the allocation proportion parameter $\hat{\pi}_{E}$ could be estimated using Thompson Sampling or derived from posterior probability-based decision rules [34].
+Similarly, the parameter $\hat{\Phi}$ can be interpreted as the posterior probability that the treatment effect exceeds zero.
+This Bayesian extension preserves the adaptive nature of the design while offering enhanced flexibility to incorporate prior information and dynamically update beliefs as data accumulate.
+An important methodological consideration pertains to the generalizability of the SAFER design.
+While the motivating example is situated within a non-inferiority framework, the design can be customised for superiority trials.
+A central message of this work is to highlight the role of safety data as a guiding component in RAR designs, owing to their early availability and their potential to enhance patient-centred research.
+Furthermore, both the $\arg\min(SAFER(\hat{\pi}_{E}))$ and $\arg\max(SAFER(\hat{\pi}_{E}))$, as well as the function boundaries (which, in our example, are defined as $\text{lower}=0.5$ and $\text{upper}=1$), can be tailored to the specific requirements of the trial setting.
+Our simulation results demonstrate that the proposed design, when simulated under a positive association between safety and efficacy, achieves power exceeding the nominal 80% planned during sample size calculation.
+This surplus is attributable to the increased probability of rejecting the null hypothesis when the experimental treatment performs better than expected.
+We demonstrate that such a power surplus can enhance the likelihood of early stopping for efficacy, thereby reducing trial duration and associated costs.
+Alternatively, this additional power could be leveraged by incorporating further adaptive elements into the design, such as sample size re-estimation or adjustment of the non-inferiority margin.
+Finally, it is worth highlighting that the clinical rationale for favouring allocation towards the arm demonstrating superior performance on the safety endpoint remains justifiable, even under the assumption of independence between endpoints and the associated loss of statistical power.
+Since the loss of power under varying allocation proportions is quantifiable, trial designs could, resources permitting, incorporate a pre-specified sample size adaptation rule at interim analyses.
+This would allow for the recovery of power losses while preserving the benefits of a safety-driven RAR design  [35].
+Limitations of the parametric assumptions
+While the exponential distribution is widely used in simulations for its simplicity and analytical convenience, it relies on the assumption of a constant hazard rate over time.
+This assumption may not always hold in real-world oncology trials, where the risk of progression or death can change due to treatment dynamics or patient-specific factors.
+For example, hazard rates may decrease initially and rise again later during follow-up.
+Although our simulations are based on this parametric model, we account for its limitations by adopting semi-parametric methods for estimation and testing, which do not rely on distributional assumptions and offer greater robustness in applied settings.
+However, the SAFER design is highly customizable: both the distributional assumptions and the type of endpoint can be relaxed.
+Without loss of generality, the proposed design can also accommodate other commonly used endpoints, such as binary outcomes, under both parametric and non-parametric frameworks.
+Conclusions
+In conclusion, our proposed safety-driven RAR design offers a significant contribution in terms of practical tools to deliver patient-centred trials by adaptive learning from quickly observable safety data to proactively mitigate the assignment to poorly tolerated treatments without negatively affecting the power of the primary trial’s objective.
+While highly relevant to Phase III oncology, its principles are broadly applicable.
+Methodological research will explore alternative distributions and non-parametric methods for small samples, with empirical validation in oncology settings essential for practical refinement.
+Finally, future efforts will focus on extending this framework to multi-arm trials, incorporating richer safety outcome data, including digital endpoints (for more frequent adaptation), and exploring the impact of non-monotonic relationship between safety and efficacy endpoints on the design characteristics.
+Author contributions
+All authors contributed to the conceptualization, methodology, and interpretation of results.
+M.V.C.
+led the development of the SAFER design, conducted the simulation study, and drafted the manuscript.
+L.P.
+and S.S.V.
+provided statistical expertise, supervised the research, and critically revised the manuscript for important intellectual content.
+All authors read and approved the final version of the manuscript.
+Acknowledgments
+The authors would like to thank David Robertson, Senior Research Associate at MRC Biostatistics Unit, and Connor Fitchett, Ph.D.
+student at MRC Biostatistics Unit, for their helpful review and insightful suggestions that improved the clarity and presentation of this work.
+The authors also acknowledge the use of large language models (LLMs) to assist with generating figures and tables, and with refining grammar and wording in the manuscript.
+The LLMs were not used for data analysis, interpretation, or original scientific writing.
+All content has been carefully reviewed and verified by the authors, who take full responsibility for the integrity and accuracy of the work.
+Financial disclosure
+The authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP), the Cusanuswerk e.V.
+(LP), and the PhD scholarship from the University of Padua (MVC).
+SSV is part of PhaseV’s advisory board.
+Conflict of interest
+The authors declare no potential conflict of interests.
+Data Availability Statement
+The code supporting the findings of this study is available upon request from the corresponding author at https://github.com/MariaVittoriaChiaruttini/SAFER-design
+Supplementary Materials
+Figure A1: Theoretical loss of power (resulting from equations 5 and 6) for the primary endpoint analysis, corresponding to an Information Fractions of 50% and allocation proportions $\pi_{E}$, from 0.5 to 0.8.
+Table A1: Overall power using exponentially distributed outcomes with fixed allocation proportion and varying information fraction at interim analysis.
+Allocation Information Fraction
 Fixed Proportion 0.2 0.3 0.4 0.5
@@
 0.80 0.64 0.63 0.63 0.63
-3.4 Simulating scenarios
-In this section, we will describe in detail the six distinct scenarios used for the simulation study.
-Scenario 1: Safety-driven RAR design without a SAFER approach: impact on power under independence between the efficacy and safety endpoints.
-This scenario assesses the impact of a safety-driven RAR design—without implementing the SAFER approach—on power, under the assumption of independence between the efficacy and safety endpoints.
-Theoretical results (in section 3.3) suggest that this impact is expected to be negative, implying a loss in power.
-Given that we set $U=15$, $\hat{\pi}_{E}$ is updated every three months until month 45, following an initial burn-in period during which a balanced randomisation is applied for the first three months.
-The median time to dose reduction or drug discontinuation is simulated at 1.5 months in the control group, while in the experimental group, it is varied across values of 1.5, 1.8, 2.25, 2.8, 3.5, 4.5, and 6 months to achieve target allocation proportions (based on the Neyman allocation) of 0.5, 0.55, 0.60, 0.65, 0.70, 0.75, and 0.80, respectively.
-Scenario 2: Safety-driven RAR design with a SAFER approach: impact on power under diverse level of dependence between the efficacy and safety endpoints.
-This scenario is designed as a counterpart to scenario 1 and aims to assess the impact of changing the allocation proportion on statistical power when the SAFER design is implemented.
-Specifically, we simulated increasing levels of association between the endpoints, with $E[\gamma_{1}]$ set to 0.001, 0.005, 0.01, 0.03, and 0.05, corresponding to very weak, weak, moderate, strong, and very strong associations, respectively.
-Under this scenario, we expect power to be preserved in the presence of weak dependence between endpoints, and to increase as dependence strengthens, potentially due to the enhanced effectiveness of the experimental treatment on the efficacy endpoint driven by prolonged well-tolerated drug exposure.
-Scenario 3: Varying elasticity of the target function in the SAFER design.
-This scenario is designed to evaluate the impact of the elastic parameter $\eta$ on the SAFER approch results.
-Specifically, scenarios 3a and 3b report the results of simulations conducted with $\eta=1$ and $\eta=5$, respectively.
-These values correspond to the absence of elasticity and a high degree of elasticity in the target SAFER function, as reported in Figure 1.
-Scenario 4: Differential timing in the observation of endpoint data.
-This scenario aims to assess the speed at which the allocation proportion reaches the target level, when the median times of the safety and efficacy endpoints are either close or far apart i.e., delayed efficacy outcome [11].
-For instance, we expect a faster pattern when the median PFS is close to the median time to dose reduction or drug discontinuation because there will be enough PFS events for the $\Phi$ parameter estimation, early in the recruitment period.
-On the other hand, when evidence of the positive association between safety and efficacy emerges later in the recruitment, we expect the SAFER function to down-weight the allocation proportion for a longer period until sufficient evidence on PFS becomes available.
-Under this scenario, the median PFS ranges from 3 to 24 months.
-Scenario 5: Impact of informative drop-out (related to the safety endpoint).
-This scenario aims to assess the impact of informative drop-out on the operating characteristics of the proposed design.
-We define informative drop-out as a situation in which a patient is observed until experiencing a side effect and subsequently drops out of the study.
-In practice, this means that the patient is followed until dose reduction or drug discontinuation occurs, after which neither the time to PFS nor the occurrence of the progression event is observed.
-We simulated the average allocation proportion and power under varying informative drop-out rates, ranging from 5% (very low) to 25% (very high), with drop-outs uniformly distributed over the accrual period.
-To address the potential bias introduced by informative drop-out, we considered the strategies for handling intercurrent events as defined in the Estimands framework [29].
-Specifically, we considered the composite strategy, where the time-to-PFS is defined as the last available follow-up time, with the occurrence of the progression = yes.
-Scenario 6: Impact of under-reported safety events.
-This scenario aims to assess the impact of under-reported safety events on the operating characteristics of the proposed design.
-Unlike the informative drop-out phenomenon—where patients are observed until an adverse event occurs and then drop out—in the case of under-reported safety events, it is not recorded at all and only the survival outcome (i.e., progression and the corresponding time-to-PFS) is available [30].
-3.5 Simulation Results
-The total sample size required to achieve at least a power value of 0.8 is 881 (about 18 patients enrolled per months).
-However, due to the interim analysis, the total number of events needed increases to 888 when IF = 0.5.
-Table 2 presents the results of the simulations under scenario 1 (independence between endpoints).
-Both the estimated allocation proportion and the observed loss of power closely align with the target allocation and the theoretical power, respectively, with only a minor discrepancy with respect to Table 1.
-This slight difference can be attributed to the burn-in period, since the results in Table 1 are derived using the unequal allocation proportion without any burn-in period, while scenario 1 includes this initial adjustment phase.
-Table 2: Estimated allocation proportion and power summary: Scenario 1
-{tblr}
-cells = c, cell11 = c=2, cell13 = c=4, vline2 = 1, vline2-6 = 2, vline3 = 3-9, hline1,10 = -0.08em, hline2-3 = -, Allocation & Information Fraction
-Target Estimated 0.2 0.3 0.4 0.5
-0.5 0.5 0.80 0.80 0.80 0.80
-0.55 0.54 0.80 0.79 0.79 0.79
-0.6 0.59 0.78 0.79 0.78 0.78
-0.65 0.64 0.77 0.77 0.76 0.76
-0.7 0.69 0.74 0.74 0.73 0.73
-0.75 0.73 0.71 0.71 0.70 0.70
-0.8 0.78 0.66 0.66 0.65 0.66
-Table 3 presents the results of the simulations under scenario 2, 3a, and 3b.
-In scenario 2, the estimated allocation proportion closely follows the target allocation, reflecting the true difference between treatment arms in terms of time to dose reduction or drug discontinuation.
-When there is no difference between arms on the safety endpoint (i.e., target allocation = 0.5), the power remains stable at approximately 0.8.
-When $E[\gamma_{1}]$ ranges from 0.001 to 0.005 (very weak or weak association between endpoints), the increasing target allocation leads to a loss of power in the efficacy endpoint analysis.
-On the contrary, when $E[\gamma_{1}]$ increases, reaching 0.03 and 0.05 values (strong or very strong positive association between endpoints), we observe a power gain.
-This improvement is driven by a biologically meaningful positive effect on PFS due to a drug exposure free from AEs (expressed by the model in section 3.2).
-Under scenario 3a, the SAFER approach is implemented with the elastic parameter $\eta$ fixed at 1.
-In the absence of endpoints association, the SAFER approach effectively shrinks the estimated allocation towards 0.5, thereby preserving statistical power.
-Unlike scenario 2, where power losses were observed, when the endpoints association is very weak or weak, in scenario 3a the power of the efficacy analysis consistently remains above 0.8, regardless of the increase in target allocations.
-Moreover, as expected, higher allocation proportions are observed for stronger association between endpoints and more extreme target allocations.
-The maximum allocation proportion achieved in scenario 3a is approximately 0.7.
-This reflects the effect of setting $\eta=1$, which enforces strong regularization and slows the adaptation of allocation proportions.
-On the other hand, in scenario 3b, when $\eta=5$, the average allocation reaches 0.74 when the target is 0.8.
-These results empirically demonstrate the role of $\eta$: the higher its value, the closer the estimated average allocation approaches the target allocation.
-Furthermore, regardless of the elasticity of the target function, when the relationship with the safety endpoint is weak, the SAFER approach consistently shrinks $\hat{\pi}_{E}$ downward.
-This mechanism minimises the potential loss of power in the efficacy analysis.
-Table 3: Estimated allocation proportion and power summary: Scenarios 2, 3a, and 3b (IF = 0.5)
-Target allocation Endpoint association Scenario 2 Scenario 3a ($\eta=1$) Scenario 3b ($\eta=5$)
-$\hat{\pi}_{E}$ (P) $SAFER(\hat{\pi}_{E})$ (P) $SAFER(\hat{\pi}_{E})$ (P)
-0.5 Very Weak 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
-Weak 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
-Moderate 0.50 (0.80) 0.50 (0.80) 0.50 (0.80)
-Strong 0.50 (0.79) 0.50 (0.79) 0.50 (0.79)
-Very Strong 0.50 (0.79) 0.50 (0.79) 0.50 (0.79)
-0.6 Very Weak 0.59 (0.78) 0.52 (0.80) 0.54 (0.80)
-Weak 0.59 (0.80) 0.53 (0.82) 0.54 (0.82)
-Moderate 0.59 (0.81) 0.53 (0.83) 0.54 (0.83)
-Strong 0.60 (0.86) 0.53 (0.87) 0.55 (0.87)
-Very Strong 0.60 (0.89) 0.53 (0.91) 0.55 (0.91)
-0.7 Very Weak 0.69 (0.74) 0.55 (0.81) 0.58 (0.81)
-Weak 0.69 (0.78) 0.55 (0.84) 0.59 (0.84)
-Moderate 0.69 (0.81) 0.56 (0.87) 0.59 (0.87)
-Strong 0.69 (0.92) 0.58 (0.95) 0.62 (0.95)
-Very Strong 0.69 (0.96) 0.60 (0.98) 0.64 (0.99)
-0.8 Very Weak 0.78 (0.67) 0.57 (0.81) 0.62 (0.81)
-Weak 0.78 (0.73) 0.59 (0.87) 0.64 (0.86)
-Moderate 0.78 (0.79) 0.60 (0.92) 0.66 (0.92)
-Strong 0.79 (0.94) 0.66 (0.99) 0.71 (0.99)
-Very Strong 0.79 (0.99) 0.70 (1.00) 0.74 (1.00)
-P = Power; $\hat{\pi}_{E}$ = estimated allocation using Neyman approach; $SAFER(\hat{\pi}_{E})$ = estimated allocation using SAFER approach.
-Figure 4 and Figure 5 illustrate the median allocation proportion, estimated using the SAFER design, over the accrual period of 48 months.
-Scenario 4 is specifically designed to evaluate the differences in the speed at which the target allocation proportion is reached.
-This is assessed under conditions where the median times to the safety and efficacy endpoints are either close together or more distant.
-As expected, a faster pattern of reaching the target is observed when the median PFS is close (3 months) to the median time to dose reduction or drug discontinuation.
-Notably, when the median PFS is 3 months, a significant reduction in variability of allocation proportion estimation is observed already from the 6th month (Figure 4).
-Moreover, when the difference in safety outcome between arm decreases i.e., the target allocation is 0.7 (Figure 5), the reduction in variability of allocation proportion estimation, in the panel A, is observed from the 12th month.
-Moreover, as the median PFS increases (panel B-D), the evidence of the positive association between the endpoints becomes more apparent later in the recruitment period.
-As a result, the SAFER elastic function consistently slows down the approach to the target allocation when efficacy and safety endpoints are temporally misaligned.
-Figure 4: Median and IQR of allocation proportions under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;).
-Target allocation=0.8.
-IF=0.5.
-$\eta$=5.
-$E[\gamma_{1}]$=0.05
-Figure 5: Median and IQR of allocation proportions under varying endpoint proximities (A: Median PFS=3 months; B: Median PFS=9 months; C: Median PFS=18 months; D: Median PFS=24 months;).
-Target allocation=0.7.
-IF=0.5.
-$\eta$=5.
-$E[\gamma_{1}]$=0.05
-Table 4 reports the average allocation proportion and the estimated power at increasing percentages of drop-out and under-reporting rates.
-A direct comparison with the values reported in Table 2 corresponding to scenario 3b and association "Very Strong" is possible since both $E[\gamma_{1}]$ and $\eta$ are fixed at the same levels.
-In scenario 5, as the drop-out rate increases, the allocation proportion also increases.
-For instance, when the target allocation is 0.8, it rises from 74% (no drop-out in Table 2) to 77% with a drop-out rate of 25%.
-Interestingly, the estimated power also increases with higher drop-out rates.
-Although this result might seem counter-intuitive, a detailed explanation is provided in the discussion section.
-On the other hand, in scenario 6, as the under-report rate increases, the allocation proportion decreases.
-This result is expected, as the core of the proposed RAR method relies on detecting differences in the safety endpoint.
-The greater the under-reporting of this endpoint, the less evidence the algorithm has to adjust the allocation proportion in favour of the better-tolerated arm.
-Table 4: Estimated allocation proportion and power summary: Scenarios 5 and 6 (IF=0.5, $E[\gamma]=0.05$, $\eta=5$)
-Target allocation Rate level Scenario 5 (Drop-out) Scenario 6 (Under-reporting)
-$SAFER(\hat{\pi}_{E})$ (P) $SAFER(\hat{\pi}_{E})$ (P)
-0.5 Very Low 0.50 (0.81) 0.50 (0.79)
-Low 0.50 (0.81) 0.50 (0.79)
-Moderate 0.50 (0.83) 0.50 (0.79)
-High 0.50 (0.85) 0.50 (0.79)
-Very High 0.50 (0.87) 0.50 (0.79)
-0.6 Very Low 0.55 (0.92) 0.55 (0.91)
-Low 0.56 (0.94) 0.54 (0.91)
-Moderate 0.56 (0.95) 0.54 (0.91)
-High 0.56 (0.96) 0.54 (0.90)
-Very High 0.56 (0.97) 0.53 (0.91)
-0.7 Very Low 0.64 (0.99) 0.63 (0.98)
-Low 0.65 (0.99) 0.62 (0.98)
-Moderate 0.65 (1.00) 0.61 (0.99)
-High 0.66 (1.00) 0.60 (0.99)
-Very High 0.67 (1.00) 0.60 (0.98)
-0.8 Very Low 0.75 (1.00) 0.73 (1.00)
-Low 0.76 (1.00) 0.72 (1.00)
-Moderate 0.76 (1.00) 0.71 (1.00)
-High 0.77 (1.00) 0.70 (1.00)
-Very High 0.77 (1.00) 0.69 (1.00)
-P = Power; $SAFER(\hat{\pi}_{E})$ = Estimated allocation using SAFER approach.
-4 Discussion
-This paper introduces a safety-driven RAR– namely SAFER design – tailored for non-inferiority trials which aims to identify alternative curative strategies—such as lower or fractionated doses, different schedules, regiments—deliver, or combo therapies with comparable efficacy to the standard of care but higher tolerability and fewer side effects.
-By incorporating safety data into the randomisation process, our method enhances patient-oriented research by preferentially allocating participants to treatment arms with a more favourable adverse event or side effect profile, while preserving statistical power.
-This approach is particularly relevant in oncology, where novel agents—often targeting specific molecular pathways—are typically administered over prolonged treatment cycles and are frequently associated with chronic toxicities, which constitute a major cause of treatment discontinuation.
-Main Results
-Our results confirm the theoretical expectation that unbalanced allocation in two-arm non-inferiority trials leads to a power loss in the efficacy endpoint analysis.
-The use of response-adaptive designs in the context of non-inferiority has been previously explored in the methodological works of [31, 32], and applied in clinical trials by [33, 34].
-These studies, however, define the study power in Bayesian perspective for multi-arm clinical trials, where the goal is to correctly identify as many non-inferior arms as possible among multiple experimental groups, under the assumption that both inferior and non-inferior treatments exist relative to a common control.
-Such approaches are not directly transferable to the more common two-arm non-inferiority trials, particularly prevalent in phase III oncology studies.
-In these settings, the alternative hypothesis explicitly posits that the experimental treatment is non-inferior to the control, and the study power is defined as the probability of correctly rejecting the null hypothesis of inferiority—i.e., demonstrating that the estimated treatment effect lies within the pre-specified non-inferiority margin.
-The trade-off between safety-driven allocation proportion and statistical power is particularly relevant in the context of oncology, where assigning more patients to better-tolerated arms mat improve patient experience and treatment effectiveness.
-The proposed SAFER approach offers a practical solution by dynamically adjusting the allocation proportion based on the observed association between safety and efficacy.
-When no association is observed, the SAFER approach appropriately shrinks the allocation back toward complete randomisation (1:1), preserving statistical power.
-When a strong (weak) positive association emerges, the algorithm allows for more (less) aggressive adaptation without undermining statistical validity.
-Moreover, the elasticity parameter ($\eta$) can regulate the speed and extent of adaptation.
-Smaller values of $\eta$ impose greater regularization, making the allocation proportion more conservative (close to 0.5), whereas higher values enable more aggressive deviation from balanced allocation.
-Moreover, we underscore the critical role of endpoint timing in the proposed design (scenario 4).
-When the safety and efficacy endpoints are temporally aligned, the design detects associations more rapidly, enabling faster and more confident adjustments in allocation proportions.
-In contrast, greater temporal separation between endpoints delays the detection of such associations and slows the adaptation process.
-This finding highlights a key advantage of the SAFER framework: the flexibility provided by the proposed function allows for better control over the degree of adaptation, particularly in settings with delayed efficacy outcomes—a common feature in survival studies.
-Unlike Sequential Maximum Likelihood Estimation RAR approach, which may adjust allocations based solely on safety endpoint data, the SAFER design incorporates regularization that prevents premature and extreme imbalances in allocation.
-This is especially relevant in non-inferiority settings, where the experimental arm could ultimately prove inferior.
-By moderating allocation shifts until sufficient evidence accumulates on the efficacy endpoint, the design protects patients from being preferentially assigned to an arm that performs better on safety outcomes but may fail on efficacy assessment.
-This guarantees that, under the null hypothesis, the allocation mechanism does not favour an inferior treatment, even if it appears more tolerable.
-Further, the study assessed the impact of two common issues in oncology trials: informative drop-out and under-reported adverse events.
-The adoption of a composite strategy to address informative drop-out results in an increase in the estimated allocation proportion as the drop-out rate rises (Table 4).
-Specifically, under this strategy, the time-to PFS is imputed as equal to the time-to-drug reduction for patients who discontinue prematurely.
-Consequently, the allocation proportion gradually converges toward $\hat{\pi}_{E}$, regardless the adaptation of the SAFER approach.
-This convergence occurs because the observed differences in PFS between treatment arms become increasingly similar to the differences in time-to-drug reduction, thereby diminishing the influence of the down-weighting effect introduced by the SAFER approach.
-In parallel, statistical power also increases.
-This is attributable to the growing number of events contributing to the composite endpoint (safety plus efficacy) as the drop-out rate rises.
-Additionally, by design, the treatment arms show a substantial difference on the safety endpoint, which enlarge the overall effect size captured by the composite measure.
-As a result, the likelihood of rejecting the null hypothesis increases.
-While a higher drop-out rate introduces bias into the estimation process, under the alternative hypothesis, this bias may not necessarily be detrimental to patient outcomes.
-However, a problematic scenario may arise if the experimental arm demonstrates better tolerability but worse efficacy.
-In such cases, the composite strategy may fail to protect the efficacy endpoint analysis.
-Specifically, imputing the time-to-PFS as equal to the time of dose reduction or discontinuation may erroneously lead to the rejection of the null hypothesis, thereby substantially increasing the risk of Type I error rate inflation.
-Importantly, this bias is not introduced by the RAR design itself, but rather by the nature of the composite strategy—particularly when drop-out is more likely in the presence of poor tolerability, and tolerability differs significantly between arms.
-Therefore, when adopting the composite strategy, the treatment effect on the efficacy endpoint should be interpreted as a combined measure of both efficacy and safety since it reflects the integrated impact of tolerability on clinical outcomes.
-Under this interpretation only, the use of a composite strategy is reasonable and justified.
-As an alternative, when a biologically plausible and well-specified association exists between the efficacy and safety endpoints, a hypothetical strategy may be used.
-Within this framework, the occurrence of the efficacy endpoint can be modeled conditionally on the observed safety outcome, thereby estimating what would have occurred in the absence of the intercurrent event [35].
-This approach can preserve estimand consistency and potentially improve inference efficiency, provided the assumptions underlying the association model are adequately met.
-On the other hand, if the efficacy data are available even for patients who drop out, a treatment policy strategy could be adopted.
-This would mitigate the risk of biased inference due to informative censoring [29, 35].
-In general, further analyses and comprehensive simulation studies are necessary to thoroughly assess the effects of informative drop-out under varying scenarios and parameter settings and the best way to handle the bias in this setting.
-Considerations on design flexibility and generalizability
-The current design adopts a group-sequential framework with a single interim analysis and a final analysis.
-However, it is fully customizable, allowing for the specification of multiple interim analyses and the selection of various $\alpha$-spending functions.
-Additionally, a $\beta$-spending function can be incorporated to formally account for futility stopping, thereby providing greater flexibility in controlling both Type I and Type II error rates [36].
-Although the proposed design is developed within a frequentist framework, a Bayesian counterpart can be readily implemented.
-Specifically, the allocation proportion parameter $\hat{\pi}_{E}$ could be estimated using Thompson Sampling or derived from posterior probability-based decision rules [37].
-Similarly, the parameter $\Phi$ can be interpreted as the posterior probability that the treatment effect exceeds zero.
-This Bayesian extension preserves the adaptive nature of the design while offering enhanced flexibility to incorporate prior information and dynamically update beliefs as data accumulate.
-An important methodological consideration pertains to the generalizability of the SAFER design.
-While the motivating example is situated within a non-inferiority framework, the design is equally applicable to superiority trials.
-A central message of this work is to highlight the role of safety data as a guiding component in RAR designs, owing to their early availability and their potential to enhance patient-centred research.
-Furthermore, both the $\arg\min(\hat{\pi}_{E}^{\text{SAFER}})$ and $\arg\max(\hat{\pi}_{E}^{\text{SAFER}})$, as well as the function boundaries (which, in our example, are defined as $\text{lower}=0.5$ and $\text{upper}=1$), can be tailored to the specific requirements of the trial setting.
-Furthermore, a transition from non-inferiority to superiority testing could be even considered within the same trial framework, especially in scenarios where safety and efficacy endpoints are positively associated.
-Our simulation results demonstrate that the design, simulated under positive association between safety and efficacy, achieves power exceeding the nominal 80% planned during sample size calculation.
-This surplus is attributable to the conservative nature of non-inferiority testing, wherein the pre-specified margin inflates the rejection region and increases the probability of rejecting the null hypothesis when the experimental treatment outperforms expectations.
-Formally, superiority can be concluded if the 95% confidence interval for the treatment effect not only excludes the non-inferiority margin but also lies entirely above zero, thereby rejecting the null hypothesis of no difference at the conventional $\alpha$ level [38].
-By reformulating the hypothesis test to directly compare mean treatment effects—i.e., $H_{0}:\theta_{E}\leq\theta_{C}$ versus $H_{1}:\theta_{E}>\theta_{C}$—and removing the non-inferiority margin from the test statistic, the same dataset can be reanalyzed for superiority.
-Notably, this procedure leverages existing data without requiring an increase in sample size.
-In favourable scenarios of positive association between endpoints, the observed power surplus could thus be reallocated to formally test for superiority, enhancing the trial’s inferential strength and its potential clinical relevance.
-Considerations on association between endpoints
-The proposed design, for simulation purposes, assumes a functional relationship between the cumulative number of completed drug cycles and the PFS rate, under the hypothesis that dose reductions/drug discontinuation adversely impact treatment efficacy.
-Specifically, the design models this association to reflect the plausible clinical scenario where reduced exposure to full-dose treatment diminishes therapeutic benefit.
-Although the motivating case study—the CAPP-IT trial—compares standard treatment versus standard treatment plus pyridoxine (experimental arm), the proposed methodological framework is generalizable to alternative settings.
-For instance, it can be applied to trials comparing two distinct treatment regimens [39], schedules [40], or dosing strategies [41], whether in monotherapy or combination settings.
-Adaptation to such contexts would simply require modification of the underlying association model linking the safety and efficacy endpoints.
-Importantly, it should be noted that the specification of the association model between safety and efficacy endpoints is required exclusively for simulation purposes to assess the operating characteristics of the SAFER design.
-In the actual trial implementation, no explicit modeling of this relationship is necessary to drive allocation decisions.
-The estimation of the $\Phi$ parameter—quantifying the empirical strength of the association—is derived directly from the same generalized linear model specified for the efficacy endpoint analysis, formally defined as $Y^{(p)}\sim j\in{E,C}$, where $E$ and $C$ denote the experimental and control arms, respectively.
-This model serves a dual purpose within the proposed adaptive framework.
-First, it is applied periodically—for instance every three months—to update the allocation proportions, where it functions solely to compute the test statistic ($z$-value) and its corresponding C.D.F.
-This C.D.F.
-is then used to inform the SAFER allocation mechanism without performing any interim hypothesis testing or inferential decision-making, avoiding uncontrolled multiple testing issues [42].
-Second, the same model underpins the formal inferential framework at both interim and final analyses, where the computed $z$-value is evaluated against pre-specified decision boundaries to support statistical testing of the primary hypothesis.
-Furthermore, it is worth highlighting that the clinical rationale for favouring allocation towards the arm demonstrating superior performance on the safety endpoint remains justifiable, even under the assumption of independence between endpoints and the associated loss of statistical power.
-Since the loss of power under varying allocation proportions is quantifiable, trial designs could, resources permitting, incorporate a pre-specified sample size adaptation rule at interim analyses.
-This would allow for the recovery of power losses while preserving the responsiveness of the randomisation mechanism [43].
-Considerations on safety endpoint analysis
-Although the present work is centred on hypothesis testing for the efficacy endpoint, the proposed design framework is sufficiently flexible to incorporate safety endpoints without loss of generality.
-These endpoints can not only be leveraged to inform the randomisation algorithm, but may also be elevated to co-primary status.
-In such cases, appropriate control of the family-wise Type I error rate can be achieved through standard procedures (e.g., Bonferroni correction, Holm’s method) [44] or more advanced alpha-spending strategies and multiplicity adjustment techniques tailored for hierarchical endpoints [45].
-Limitations of the parametric assumptions
-Although the exponential distribution is commonly used for simulating survival trial designs, it introduces limitations due to the parametric assumptions it imposes.
-However, the SAFER design is highly customizable: both the distributional assumptions and the type of endpoint can be relaxed.
-In the current work, we adopted time-to-event endpoints, but without loss of generality, the proposed design can also accommodate other commonly used endpoints, such as binary or count outcomes, under both parametric and non-parametric frameworks.
-Conclusions
-Ultimately, our proposed safety-driven RAR design offers a significant contribution in terms of practical tools to deliver patient-centred trials by adaptive learning from quickly observable safety data to proactively mitigate the assignment to poorly tolerated treatments without negatively affecting the power of the primary trial’s objective.
-While highly relevant to Phase III oncology, its principles are broadly applicable.
-Future efforts will focus on extending this framework to multi-arm trials and incorporating richer safety outcome data, including digital endpoints for more frequent adaptation.
-Methodological research will explore alternative distributions and non-parametric methods for small samples, with empirical validation in oncology settings essential for practical refinement.
-Acknowledgments
-The authors acknowledge the use of large language models (LLMs) to assist with generating figures and tables and refining grammar and wording in this paper.
-The LLMs were not used for data analysis, interpretation, or original scientific writing.
-All content has been carefully reviewed and verified by the authors, who take full responsibility for the integrity and accuracy of the work.
-Moreover, the authors acknowledge funding and support from the UK Medical Research Council (grants MC UU 00002/15 and MC UU 00040/03 (SSV), as well as an MRC Biostatistics Unit Core Studentship (LP), the Cusanuswerk e.V.
-(LP), and the PhD scholarship from the University of Padua (MVC).
-SSV is part of PhaseV’s advisory board.